### v2

# C2P: Featuring Large Language Models with Causal Reasoning

Abstract
Causal reasoning is the primary bottleneck that Large Language Models (LLMs) must overcome to attain human-level intelligence. To address this, we introduce the Causal Chain of Prompting (C2P) as the first reasoning framework that equips current LLMs with causal reasoning capabilities. C2P operates autonomously, avoiding reliance on external tools or modules during both the causal learning and reasoning phases, and can be seamlessly implemented during the training or fine-tuning of LLMs. Experimental results across various benchmark datasets demonstrate a significant improvement in causal learning and subsequent reasoning accuracy of LLMs. We illustrate how C2P enhances LLMs’ ability to causally reason in real-world scenarios, addressing complex problems in fields such as healthcare, medicine, economics, education, social sciences, environmental science, and marketing. With few-shot learning, GPT-4 Turbo using C2P with as few as six examples achieves significant performance improvements, boasting over a $33\%$ increase in reasoning accuracy over the most state-of-the-art LLMs, which perform nearly randomly in similar circumstances. This demonstrates the transformative potential of integrating C2P into LLM training or fine-tuning processes, thereby empowering these models with advanced causal reasoning capabilities.
1 Introduction
Recent advancements in Large Language Models (LLMs) have impacted existing AI paradigms and heightened expectations regarding AI’s capabilities (Achiam et al., 2023; Brown et al., 2020).
Figure 1: Example of the standard prompting vs few-shot learned GPT-4 with C2P in open problem in astrophysics (Pasquato et al., 2023)
Despite significant architectural differences in LLMs, they generally produce outputs based on the most likely results learned from vast amounts of training data (Vaswani et al., 2017). This enables them to acquire extensive knowledge ranging from common sense to specialized domains such as mathematics and science (Jiralerspong et al., 2024). Despite this, the inefficiency of LLMs in addressing causal reasoning questions remains their primary bottleneck and simple tasks can completely break down reasoning in state-of-the-art LLMs (Nezhurina et al., 2024). Additionally, studies such as Kalai & Vempala (2023); Xu et al. (2024) have demonstrated that, despite the training data containing numerous examples of interventions, outcomes, and explanations, as well as similar tasks, hallucinatory responses persist and there is a lack of causal reasoning capability. As a result, while they may talk causality, they are not causal (Zečević et al., 2023). This deficiency represents a fundamental drawback of LLMs as AI systems compared to human intelligence, which goes beyond mere correlations and depends on causal relationships for decision-making (Penn & Povinelli, 2007; Anwar et al., 2024).
Recently, answering cause and effect questions with LLMs has gained extensive interest (Shin et al., 2020; Jin et al., 2023b; Ashwani et al., 2024). To address this issue, it is important to note that LLMs have been utilized in conjunction with external tools to extract causal structures, as demonstrated in (Jiralerspong et al., 2024). However, their architectures lack specialized modules specifically designed to enhance the understanding of cause-and-effect relationships within their outputs (Wang et al., 2023; Imani et al., 2023). Aside from studies that reason causally based on the knowledge that is already in their training data (Petroni et al., 2019; Jiang et al., 2020) and the ones that use LLMs in causality (Kıcıman et al., 2023; Zhang et al., 2023; Feder et al., 2024; Khatibi et al., 2024), chain-of-thought prompting is presented in (Wei et al., 2022) as one of the initial attempts in enhancing the reasoning in LLMs that shows improvement based on the data of the given query. However, LLMs still struggle with rigorous numerical and abstract reasoning among many other tasks (Xu et al., 2023). For example, a recent work, Causal Reasoning Assessment Benchmark (CRAB, Romanou et al., 2023), is designed to evaluate the causal understanding of events in real-world narratives. This study demonstrated that most systems perform poorly in identifying cause-and-effect tasks. Similarly, in (Jin et al., 2023b), the CORR2CAUSE dataset is introduced and demonstrated that current models often perform no better than random chance when tasked with causal questions. Following that, in (Jin et al., 2023a), the CLADDER dataset is introduced to asses Average Treatment Effects with LLMs and it is demonstrated that these models struggle with causal tasks. In their study, with the implementation of the proposed framework, CAUSALCoT, progress has been made in evaluating the average treatment effect where LLMs are provided with a collection of causal graphs and various types of queries (associational, interventional, and counterfactual), such as those included in the CLADDER dataset. More recently, in (Ashwani et al., 2024), a novel architecture called the Context-Aware Reasoning Enhancement with Counterfactual Analysis (CARE-CA) framework is presented to enhance causal reasoning and explainability. Their proposed framework incorporates an external explicit causal detection module with ConceptNet (Speer et al., 2017) and counterfactual statements, as well as implicit causal detection through LLMs, showing progress in causal reasoning in short and simple queries. Several other works at the intersection of causal inference and LLMs are discussed in an extensive survey by Liu et al. (2024).
As argued by Pearl (1995), causal Directed Acyclic Graphs (DAGs), along with d-separation, allow for the investigation of cause-and-effect relationships without relying on structural equation models in computational studies. Inspired by Pearl’s foundational work, we propose a novel framework, named Causal Chain of Prompting (C2P), to address the inefficiencies of LLMs in handling causal queries (see Figure 1 as an example). We show that by identifying the adjacency matrix, equivalent to and instead of the causal DAG in Pearl’s work, the causal relationships among the variables in the premise’s cause-and-effect relationships can be effectively reasoned within the context of language models. Contrary to the existing research in addressing weaknesses of LLMs in causal reasoning, the C2P operates autonomously, avoiding reliance on external tools or modules during both the learning and reasoning phases to answer causal questions. Additionally, C2P can easily be implemented in the training or fine-tuning process of LLMs. C2P includes five simple main sub-tasks, as follows: (1) Prompting for extracting the random variables from the provided data. (2) Prompting for extracting all conditional and unconditional relations and cause and effect relations specifically mentioned among the random variables. (3) Prompting to create the initial adjacency matrix with values 1 for all elements except the diagonal elements and the ones corresponding to effect-cause relations (the cause and effects elements are also one). (4) Prompting of the conditional and unconditional independencies and identification of the colliders, step by step, to extract the causal adjacency matrix; and, (5) Prompting for reasoning questions or hypotheses. To evaluate the accuracy and reliability of implementing the C2P on LLMs, we initially assess it using publicly available synthetic datasets such as (Jin et al., 2023b). Subsequently, we evaluate it in more realistic and complex scenarios in real-world problems presented in (Pearl & Mackenzie, 2018) and (Pasquato et al., 2023). Moreover, we present results on few-shot learning with C2P on both synthetics and realistic scenarios.
Contributions.
In this work, we present several important contributions to facilitate causal reasoning in language models. Concretely,
  1. 1.
    We introduce the C2P framework as the first reasoning framework to equip LLMs with the ability of causal reasoning within real-world scenarios, without relying on external tools.
  2. 2.
    Through extensive experiments with our framework, we demonstrate a significant improvement of LLMs in causal reasoning in various benchmarks. Additionally, we examine the performance of the C2P framework on more complex and real-world scenarios in various domains.
  3. 3.
    We perform few-shot learning experiments in section 4 with GPT-4 Turbo using our C2P framework and show how integrating C2P during the training or fine-tuning of LLMs can revolutionize existing models. This approach equips them with causal reasoning capabilities akin to the transformative impact of ‘Chain of Tought’ (Wei et al., 2022) on LLMs, as discussed by Chung et al. (2024).
The codes for sections 3 and 4 are publicly available at https://anonymous.4open.science/r/C2P-5C2A/.
2 Preliminaries on Causal Learning and Reasoning
To lay the groundwork for our framework, understanding the fundamental concept of cause and effect is paramount. A foundational and widely acknowledged principle in causality is the one proposed by Reichenbach, which posits the following:
Principle 1 (Common cause principle (Reichenbach, 1991))
[…]
Markov Property: The Markov property in a DAG $G$ states that each node $X_{i}$ is conditionally independent of its non-descendants given its parents, namely $X_{i}\perp\!\!\!\perp\text{NonDe}(X_{i})|\text{Pa}(X_{i})$, where $\text{NonDe}(X_{i})$ denotes the non-descendants of $X_{i}$ excluding itself, and $\text{Pa}(X_{i})$ denotes the parents of $X_{i}$. Using the Markov property, we can factorize the joint distribution of all the nodes in the graph into $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i}|\text{Pa}(X_{i}))$. To infer the causal graph from probability distributions, a common assumption is faithfulness, namely the validity to infer all the d-separation sets in the graph from the independence relations in the probability distribution. In our work, we also take this broadly taken assumption which holds for most real-world scenarios.
Directed Acyclic graph: A graph $G$ is called a partially directed acyclic graph (PDAG) if there is no directed cycle, that is, if there is no pair $(j,k)$ with directed paths from $j$ to $k$ and from k to $j$. $G$ is called a directed acyclic graph (DAG) if it is a PDAG and all edges are directed.
Markov Equivalence of Graphs: We denote two DAGs as Markov equivalent if they induce the same joint distribution $P(\mathbf{X})$. The set of DAGs that are Markov equivalent to each other is called a Markov equivalence class (MEC). Causal graphs in the same MEC can be easily identified since they have the same skeleton (i.e., undirected edges) and V-structures (i.e., structures in the form of $X_{1}\rightarrow X_{2}\leftarrow X_{3}$ where $X_{1}$ and $X_{3}$ are not connected).
The be able to perform any level of causal reasoning, the Ladder of Causation is proposed with three main rungs, “seeing”, “doing”, and “imagining”, in (Pearl & Mackenzie, 2018). To fulfill all rungs of the ladder, more information on the causal structure is needed, which can be mostly provided with structural causal models (SCM) (Assuming all the assumptions are satisfied) (Bareinboim et al., 2022). The current causal discovery methods are divided into two groups, constraint-based algorithm such as the PC algorithm, which has a PDAG as an output, that represents the MEC of the true underlying graph and is the best result such methods can extract. Second is score-based methods such as NOTEARS (Zheng et al., 2018), GOLEM (Ng et al., 2020), and DAGMA (Bello et al., 2022), among many others, that extract a DAG that mostly fits the data and demonstrated very high accuracy in extracting Bayesian networks. Score-based methods require solving an optimization problem based on the numerical values of variables, rendering them impractical for implementation in LLM models.
PC algorithm: Following these materials, the PC algorithm is developed based on Reichenbach’s common cause principle and Markov property as follows for a causal mechanism with $p$ random variables (Spirtes et al., 2001);
  1. (i)
    Form a complete undirected graph
  2. (ii)
    Eliminate edges between variables that are unconditionally independent
  3. (iii)
    For each pair of variables $(X_{1},X_{2})$ having an edge between them, and for each variable, $X_{3}$ with an edge connected to either of them, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|X_{3}$
  4. (iv)
    For each pair of variables $X_{1},X_{2}$ having an edge between them, and for each pair of variables $\{X_{3},X_{4}\}$ with edges both connected to $X_{1}$ or both connected to $B$, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|\{X_{3},X_{4}\}$.
  5. (v)
    For each triple of variables $(X_{1},X_{2},X_{3})$ such that $X_{1}$ and $X_{2}$ are adjacent, $X_{2}$ and $X_{3}$ are adjacent, and $A$ and $C$ are not adjacent, orient the edges $X_{1}-X_{2}-X_{3}$ as $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, if $X_{2}$ was not in the set conditioning on which $X_{1}$ and $X_{3}$ became independent and the edge between them was accordingly eliminated. We call such a triple of variables a V-structure.
  6. (vi)
    For each triple of variables such that $X_{1}\rightarrow X_{2}-X_{3}$, and $X_{1}$ and $X_{2}$ are not adjacent, orient the edge $X_{2}-X_{3}$ as $X_{2}\rightarrow X_{3}$. This is called orientation propagation.
d-separation and do calculus (Pearl, 1995): do-calculus, developed by Judea Pearl and others, is a set of rules used to transform and manipulate causal expressions within causal diagrams (or graphical models). It is a generalization of do-calculus, which is itself a formal tool used to reason in causal relationships from a mixture of experimental and observational data. do-calculus consists of three main rules that enable one to rewrite expressions involving interventions (typically represented as $do(x)$, indicating an intervention to set variable $X$ to value $x$):
  1. 1.
    Insertion/deletion of observations: If a variable $X_{2}$ is independent of $X_{1}$ given a set of other variables $X_{3}$, then the conditional distribution of $X_{2}$ given $do(X_{1})$ and $X_{3}$ is the same as the conditional distribution of $X_{2}$ given $X_{3}$ alone. This rule can be applied when $X_{1}$ and $X_{2}$ are d-separated by $X_{3}$ in the graph modified by removing all arrows pointing into $X_{1}$.
  2. 2.
    Action/observation exchange: If $X_{2}$ is independent of $X_{1}$ given $X_{3}$ and another set $W$ where $X_{1}$ is intervened, then the graph can be modified to reflect this exchange.
  3. 3.
    Insertion/deletion of actions: This rule states that if intervening on $X_{1}$ does not affect $X_{2}$ after accounting for $X_{3}$, the intervention can be ignored in the expression.
These rules are crucial for determining the identifiability of causal effects from data, allowing researchers to reason in causal relationships from a combination of experimental and observational data.
d-Separation is a criterion used in the analysis of Bayesian networks to determine whether a set of variables $X_{1}$ is independent of another set of variables $X_{2}$, given a third set of variables $X_{3}$. This concept is foundational in understanding the flow of causal influence in graphical models and helps in deciding whether a path between two variables is "blocked" or not by conditioning on other variables. The basic rule of d-separation is: A path between two variables is blocked if it includes an intermediate variable that is a collider and is not conditioned on or a non-collider that is conditioned on. Here, a collider is a variable that has inward arrows from two other nodes (i.e., $X_{1}\rightarrow X_{2}\leftarrow X_{3}$), whereas a non-collider does not meet this criterion.
3 Developing Causal Chain of Prompting
To develop C2P framework, we aim to extract the adjacency matrix of variables in a given premise. The adjacency matrix acts as equivalent to and instead of the causal DAG in Pearl’s work leading to answering the causal reasoning question. In developing the Causal Chain of Prompting, we use a few steps similar to the PC algorithm for the given premise.
3.1 Causal Chain of Prompting (C2P)
The C2P frameworks consist of 5 main sub-tasks as follows, for learning and reasoning cause and effect relation for the given premise;
  1. 1.
    Prompting to extract the random variables in the provided data.
  2. 2.
    Prompting to extract all the cause and effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise.
  3. 3.
    Prompting to create an initial adjacency matrix where all elements are 1, except for the diagonal elements and the elements corresponding to the cause-and-effect relationships specifically mentioned in the given premise (extracted in sub-task 2).
  4. 4.
    Prompting of the conditional and unconditional independency valuation and identifying the colliders, step by step, to extract the causal PDAG.
  5. 5.
    Prompting for cause-and-effect questions or hypotheses.
Each sub-task in C2P includes multiple steps and overall, to execute the frameworks, 9 main prompts have to be completed. The exact prompts for all the steps are provided in the Appendix, A.1.
3.2 Few-shot learning agenda
LLMs have powerful zero-shot capabilities, yet they might struggle with complex tasks with insufficient examples in the training process. In these situations, few-shot learning is a versatile and efficient technique for in-context learning that can be used to quickly adapt LLMs to new tasks and significantly enhance their performance (Min et al., 2022; Touvron et al., 2023). This approach involves providing several examples with desired answers, which condition the LLMs to produce the correct response for new instances with similar patterns. The few-shot learning process of C2P is similar to the previous subsection, with the prompts being based on the more abstract prompts presented there (Due to the token limitation). The prompts and an example of the given story are included in the Appendix, A.2. Depending on the token limitation of employed LLMs, the number of given examples (shots) varies. Comparing the token limitation of GPTs, we use GPT-4 Turbo, which has a 30000 token limit.
Steps from other constraint-based methods can be used instead or along with some steps of the PC algorithm in C2P methods, comprehensively discussed in (Glymour et al., 2019). For instance, steps in the FCI method can be used in cases where the causal sufficiency assumption is violated, i.e., latent variables and selection bias may be present (Spirtes et al., 1995). Additionally, it is important that the classifiers proposed by Ceraolo et al. (2024) that formalize the definition of causal questions and establish a taxonomy for finer-grained classification can be used before our framework to identify the causal question as a prerequisite of our method.
4 Results of the experiment for featuring LLMs with Causal Reasoning
4.1 Experimental setup
The experimental setup is divided into two data groups: synthetic dataset and realistic scenarios.
In a first synthetic dataset, and similar to the experimental setup of (Jin et al., 2023b), we set up the following list of LLMs for the experiments on the CORR2CAUSE dataset. Various types of LLMs are tested in (Jin et al., 2023b) and we only include powerful LLMs in our experiment to avoid reproducing the same results. To test existing LLMs, we first include two BERT-based NLI models in the transformers library (Wolf et al., 2020): BART (Lewis et al., 2019), DistilBART (Shleifer & Rush, 2020). We also evaluate the recent, more efficient models, LLaMa3-8B (Touvron et al., 2023) and LLaMa3-70B (Taori et al., 2023). We assess the GPT-3.5 (i.e., ChatGPT), and the latest GPT-4 (OpenAI, 2023) by April 2023, using the OpenAI API (https://openai.com/api/) with temperature 0. We employed the most recent GPTO as well. Then, we assess C2P on a proposed synthetic dataset, similar to the one in (Jin et al., 2023a), but more complex and with less information provided on the premises. Then we test the reliability of C2P in different numbers of nodes. Due to the limitation of tokens for different versions of GPTs, we employed GP4-Turbo which has a 30000 maximum token limit. We selected 100 samples at random from the corr2cause dataset for our experiment, ensuring a balanced distribution with 50 "Yes" and 50 "No" answers. In the corr2cause dataset, the majority of responses are "No", as a result, they used the F1 score as the primary metric for accuracy. Our experiment design enhances both the realism and comprehensibility of the results. These samples explore reasoning within three causal scenarios: direct cause-and-effect relationships (Fig. 2 i), indirect cause-and-effect relationships (Fig. 2 ii), and the presence of an effect due to two causes (Fig. 2 iii).
Figure 2: All possible in cause and effect relations. i. $X_{1}$ directly causes $X_{2}$ ($X_{2}$ is directly effect of $X_{1}$). ii. $X_{1}$ indirectly causes $X_{3}$ ($X_{3}$ is indirectly effect of $X_{1}$). iii. $X_{1}$ and $X_{3}$ are causes of $X_{2}$ ($X_{2}$ is common effect of $X_{1}$ and $X_{3}$)
To evaluate the proposed framework on natural stories, we utilized the state-of-the-art large language model, GPT-4, which excels at story generation. We crafted detailed instructions in the prompts and generated approximately 20 stories for our case study in fields such as healthcare and medicine, economics, education, social sciences, environmental science, and marketing, all emphasizing the importance of causality. This process is similar to the one presented in (Jin et al., 2023b), but our generated stories are more complex. Our approach can be tested using the examples of the Simpson’s Paradox discussed in (Pearl & Mackenzie, 2018). However, since these examples are already included in the training data of current LLMs, the models simply repeat the correct answers based on that training data, similar to the parrot study in (Zečević et al., 2023). Consequently, our generated natural stories replicate these examples in a manner that the LLM cannot address within its existing training data.
Additionally, we assess the coevolution of supermassive black holes (SMBHs) and their host galaxies using our proposed framework, replicating the results provided in (Pasquato et al., 2023).
4.2 Evaluation of the C2P on synthetic dataset
Results of the C2P on CORR2CAUSE dataset: In Table LABEL:table4, we show the performance of the more powerful LLMs in the cause-effect task reported in (Jin et al., 2023b) compared to the results of the C2P framework.
Table 1: Main results: Comparison of using C2P frameworks compared to the existing LLMs
Models F1 Precision Recall Accuracy
Random Baselines
Random (Proportional) 13.5 12.53 14.62 71.46
Random (Uniform) 20.38 15.11 31.29 62.78
BERT-Based Models
DistilBART MNLI 26.74 15.92 83.63 30.23
BART MNLI 33.38 31.59 35.38 78.50
LLaMa-Based Models
LLaMa3-8B 48.97 51.06 47.05 51
LLaMa3-70B 70.8 60.7 56.6 60
GPT-Based Models
GPT-3.5 48.8 55 44 54
GPT-4 48.4 53.3 48 53
GPT-4O 57.7 59.5 56 59
C2P on GPT-4 Turbo 92.78 95.74 90 93
Few-shot learned GPT-4 Turbo 81.06 82.97 79.24 81
According to Table LABEL:table4, performing pure causal reasoning tasks presents a significant challenge for all existing LLMs. Notably, several models perform worse than a random guess, indicating a complete failure in this specific task of pure causal reason. Nevertheless, applying our proposed framework to GPT-4 Turbo has boosted at least 33% in the model’s ability to respond to causal inquiries, suggesting a substantial improvement over the existing capabilities of LLMs. It is important to note that the primary factor contributing to the discrepancy between the LLaMa and GPT results and those reported in (Jin et al., 2023b) stems from the distribution of the provided premise. This table indicates that when reasoning questions are posed to LLMs, the responses slightly tend towards the "No" answer.
Robustness analysis: To assess the robustness of the proposed prompts, Table 2 shows the computed accuracy computed in each sub-task of implementing C2P on GPT-4 for different numbers of variables in the given premise.
Table 2: Accuracy by Number of Variables and sub-tasks
Number of variables Accuracy
First sub-task Second sub-task Third sub-task Fourth sub-task Fifth sub-task
3 variables 100% 100% 100% 100% 99%
4 variables 100% 100% 100% 98% 83.2%
5 variables 100% 100% 100% 88.3% 75.2%
6 variables 100% 100% 100% 79.1% 70.1%
4.3 Evaluation of the C2P on natural stories
To evaluate the proposed framework on natural stories, we utilize the same process presented in (Jin et al., 2023b) and generate premises. The result of applying the C2P framework, both step-by-step and few-shot learned, is provided in Table 3 and compared to the result of GPTs. This table clearly demonstrates that the existing method responds randomly to reasoning prompts. It is important to note that even when GPTs provide the correct final answer, the reasoning process leading to that answer is still incorrect. In contrast, GPT-4 Turbo, enhanced with step-by-step prompting of the C2P framework and few-shot learned C2P, boosts the accuracy of these models in similar scenarios involving natural stories.
Table 3: Result of applying the C2P framework and GPTs, TP: True Positives, FP: False Positives, TN: true negatives, FN: False Negatives
Models TP FP TN FN
GPTs
GPT-4.5 2 2 3 3
GPT-4 2 2 3 3
GPT-4o 3 2 3 3
Featured GPT
C2P on GPT-4 Turbo 5 0 5 0
Few-shot learned GPT-4 Turbo 4 0 5 1
Additionally, Table 4 provides an example of such natural stories and demonstrates the result of applying a few-shot learned GP4-Turbo with C2P to it. This showcases how GPT models (including versions 3.5, 4, 4 Turbo, and 4o) fail to respond correctly. It is important to note that the few-shot learning process of GPT-4 Turbo was performed using only six examples from the CORR2CAUSE dataset, without including any natural story examples. Note that there are no natural stories employed in the few-shot learning process.
Table 4: Example of a natural story and the result of few-shot learned GPT-4-Turbo with C2P
Example of true story Results of few-shot learned GPT-4 Turbo with C2P
Premise: eating junk food (A), obesity (C), and watching television (B) have relations with each other. There is a correlation between eating junk food and obesity, and between watching television and obesity. However, eating junk food and watching television are independent from each other. Hypothesis: Eating junk food directly affects obesity Step 1: Number of random variable:3 Names of random variable: A, B, C
Step 2: All of Statistical Relations:{ Dependencies: [[A, C], [B, C]], Unconditional Independencies: [[A, B]], Conditional Independencies: []}
Step 3: A: {A: 0, B: 1, C: 1}, B: {A: 1, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
Step 4: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
Step 5: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
Step 6: C: [A, B]
Step 7: C: [A, B]
Step 8: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 0, B: 0, C: 0}
Step 9: Checking matrix[A][D] = 1 and matrix[D][A] = 0. According to rule 2, this suggests A is a direct cause of D, or D is a direct effect of A.
4.4 Evaluation of the C2P framework on the (co)evolution of supermassive black holes and their host galaxies
In astrophysics, experiments are impossible. Thus, causal reasoning must rely exclusively on observational data. Similar to the study in (Pasquato et al., 2023), we (co)evolute supermassive black holes (SMBHs) and their host galaxies. We implement C2P on GPT-4 Turbo to a comprehensive catalog of galaxy properties to obtain a PDAG and then reason the causal questions.
Fig. 3 shows the results of each sub-tasks of implementing C2P
Figure 3: Prompts (Q) and results (A) of sub-tasks application of the C2P framework to real-world complex scenarios and steps of sub-task 3 for the given premise.
According to the extracted PDAG, many questions such as the causal given can be answered which is too complicated for the existing LLMs to answer. It is important to note that by implementing C2P, not only many causal questions can be answered with this approach, but it is easy to show which causal questions can not be answered based on the given premise and which more information has to be given to be able to causally reason. More importantly, it is worth noting that this is not the limitation of the C2P framework, any rational agent needs such information to answer such causal questions.
5 Discussion, Limitations and Future Work
Causal reasoning and identifiability: Having a PDAG of an SCM of a causal mechanism is the most effective way to achieve this with observational data. Based on this, two main questions arise: Given a PDAG, under what conditions can we make causal reasoning? This involves determining the necessary assumptions and data required to address a causal question—this issue is known as identification. It is generally possible to answer some cause-effect questions based on the PDAG, represented by a causal graph, and lower-level data. These questions are mainly the ones that are related to the part where we have extracted a collider, or if the question is just about the existence of a cause-effect relation; in some cases, these questions can be answered as well. However, it is almost impossible to answer all the causal questions based on the given information, as discussed in (Bareinboim et al., 2022; Pearl & Bareinboim, 2022). As a result, the second question is, what else is needed to answer all the causal questions? In such cases, interventional data is necessary to explicitly extract the direction of causal relations in a mechanism. To do so, while a naive baseline approach would require $O(n^{2})$ interventions, various methods have been proposed, such as those in (Kocaoglu et al., 2017; Choo & Shiragur, 2023; von Kügelgen et al., 2024; Squires & Uhler, 2023), for cases in different situations. However, one clear thing is that it is not possible to reason all the causal questions. For instance,
  • Directions cannot be discovered: PDAGs typically include both directed and undirected edges. The undirected edges indicate uncertainty about the direction of causality. It is important to note that without further data or assumptions, it is not possible to definitively determine the causal direction for these relationships with any method.
  • Full Causal Path Analysis: While PDAGs can indicate possible paths between variables, they may not fully reveal which paths are indeed causal and which are due to confounding or indirect effects. Questions about specific causal pathways can thus be hard to answer definitively.
  • Predictions Under Interventions: Questions about the outcomes of hypothetical or actual interventions on one or more variables (do-calculus questions) often require a fully specified causal model. PDAGs, with their partial specification, might not support detailed predictions under interventions without resolving the ambiguities in causal direction.
Causal reasoning is a broad and complex task in AI and LLMs. While current machine learning methods struggle with extracting causal structures and subsequently reasoning causally, this problem is even more intricate in the context of LLMs. The primary reason for this complexity is the inherent structure of LLMs and their reliance on the concept of attention. To enable LLMs to reason causally, the first step is to identify the causal question. Recently, a comprehensive database was compiled, and an AI model was trained to identify causal questions (Ceraolo et al., 2024). This model can be used prior to applying our framework to determine whether a causal reasoning process has to be activated or not. Additionally, more comprehensive examples and scenarios need to be generated to aid in the learning process of an LLM. Due to the vast variety of causal scenarios and the token limitations of current models, our approach must be integrated into the learning or fine-tuning process of an LLM. This ensures that the model can effectively learn and apply causal reasoning within its capabilities.
Scaling Instruction-Finetuned Language Models: LLMs primarily function through pattern recognition without the capability to recognize causal relationships, as indicated by prior research . This limitation highlights the importance of implementing these models with causal analysis techniques to enhance their causal reasoning ability. Previous attempts to boost LLMs’ causal reasoning involved using external tools or manually crafted prompts, which, although insightful, are impractical for easy integration. Our study introduced an algorithm that leverages LLMs’ inherent abilities to foster causal reasoning autonomously. Our findings demonstrate a marked increase in the accuracy of LLMs’ causal understanding, even in scenarios involving complex causal interactions. Additionally, incorporating a chain of thought into LLMs’ training has been proven to enhance performance across various tasks (Chung et al., 2024). Based on the extracted results on the benchmark and provided datasets in our study, and the results of the few-shot learning process, we strongly hypothesize that our approach could similarly benefit the training of language models, particularly in boosting their causal reasoning, even in smaller models that would lead to a revelation in LLMs. This enhancement could be realized using a carefully selected set of examples or synthetic samples created by LLMs, following the guidelines outlined in our algorithm.
Future works As demonstrated in the results section, C2P has significantly enhanced the reasoning capabilities of LLMs. We have shown that extracting the adjacency matrix of the variables in a given premise can function as the DAG in Pearl’s works, leading to improved reasoning ability. Our experimental results of few-shot learned LLMs indicate a significant increase in accuracy for tasks where users inquire about direct cause-and-effect relationships or common cause questions. However, the models still struggle with indirect cause-and-effect questions. Pearl demonstrated that the concepts of a DAG and d-separation complement each other in causal reasoning tasks. To address the low accuracy in answering indirect causal reasoning questions, more sub-tasks can be introduced to LLMs to perform d-separation in a more structured manner. As mentioned, due to the token limitation of LLMs, we were able to use only six or even fewer examples in the few-shot process of GPT-4-Turbo. As a result, our proposed framework can be seamlessly integrated into the fine-tuning or learning process of an LLM by introducing more extensive examples with more complex scenarios. Given the accuracy achieved in few-shot learned LLMs, the fine-tuned or trained LLMs with our approach can achieve highly accurate results, effectively addressing the causal reasoning bottleneck of LLMs.
Appendix A Appendix
A.1 Prompts for C2P frameworks
The proposed prompts for step-by-step proceeding with C2P are as follows:
  1. 1.
    Prompting for extracting the random variables from the provided data.
    1. (i)
      Please give me the number of random variables in the given premise
    2. (ii)
      Write the names of all random variables.
  2. 2.
    Prompting to extract all the cause and effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise
    • If 2 random variables, for instance, $R_{1}$ and $R_{2}$, are independent, write it in this form: "$R_{1}$ is independent of $R_{2}$". If there exist 2 random variables, for instance, $R_{1}$ and $R_{2}$, are conditionally independent given a third random variable, for instance, $R_{3}$, write it in this form: "$R_{1}$ and $R_{2}$ are independent given $R_{3}$". If two random variables, for instance, $R_{1}$ and $R_{2}$, are specially mentioned to have cause and effect relation, write it in this form: "$R_{1}$ is the cause of $R_{2}$".
  3. 3.
    Prompting to create an adjacency matrix where all elements are 1, except for the diagonal elements and the elements corresponding to the cause-and-effect relationships specifically mentioned in the given premise.
    • In this phase, each random variable is treated as a node within a fully connected undirected graph. Then, for each pair, for instance, $R_{1}$ and $R_{2}$, presented in the form: "$R_{1}$ is the cause of $R_{2}$", set the element in ["$R_{2}$", "$R_{1}$"] in the adjacency matrix to 0.
  4. 4.
    Prompting of the conditional and unconditional independency valuation and identifying the colliders, step by step, to extract the causal PDAG.
    1. (i)
      Update the adjacency matrix based on the specified unconditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent, set their corresponding entries in the adjacency matrix to 0.
    2. (ii)
      Update the adjacency matrix based on the specified conditional independencies between random variables. Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix. The initial adjacency matrix and the list of independencies are provided below. Please ensure all independencies are correctly reflected in the updated matrix. Instructions: - For each pair of variables listed as independent given other variable(s), set their corresponding entries in the adjacency matrix to 0.
    3. (iii)
      Task: Given an initial adjacency matrix, follow these steps: Step 1: Identify all rows (key values) in the matrix where there are two or more than two columns with the value "1" in them. For each identified row, find all pairs of different columns where the values are "1".Ensure to exclude rows that do not contain any pairs from the results. Step 2: Display these pairs, "All Pairs", where each row name is key, and the value is a list of column names that are identified in Step 1.
    4. (iv)
      Given the "All Pairs" and the list of independencies, follow these instructions step by step: Instruction: For each key in "All Pairs", delete all the pairs that are not mentioned as independent in the "independencies" list and return other with all their values. The "All Pairs" contains pairs of elements associated with each key. The goal is to update this by removing pairs that are not mentioned as independent. The list of independencies provides information about which pairs are independent of each other.
    5. (v)
      Given the initial adjacency matrix represented and the "All Pairs" list, for each key-value pair ("$R$") in "All Pairs", modify the initial adjacency matrix as follows: 1- Set the value in the "$C_{1}$" row and "$R$" column to 0: ("$C_{1}$", "$R$") = 0. 2- Set the value in the "$C_{2}$" row and "$R$" column to 0: ("$C_{2}$", "$R$") = 0.
  5. 5.
    Prompting for cause-and-effect questions or hypotheses
    • To extract and understand causal relations in the adjacency matrix: For each specified variable "$R$" and "$C$", for instance, that are listed in the adjacency matrix: - If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 1, then the causal direction between "$R$" and "$C$" is undetermined. - If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 0, then "$R$" is a direct cause of "$C$" or "$C$" is a direct effect of "$R$". - If matrix entry at ["$R$", "$C$"] = 0 and ["$C$", "$R$"] = 1, then "$C$" is a direct cause of "$R$" or "$R$" is a direct effect of "$C$". If two variables directly affect a third variable, the first two variables are common causes, and the third variable is a collider. Evaluate the hypothesis based on the given partially directed acyclic graph presented as an adjacency matrix with the given Instruction.
A.2 Prompts for few-shot learning of C2P
An example of given prompts for the few-shot learning process of C2P is as follows:
Premise: Suppose there is a closed system of 5 variables, A, B, C, D, and E. All the statistical relations among these 5 variables are as follows: A correlates with C. A correlates with D. A correlates with E. B correlates with D. B correlates with E. C correlates with D. C correlates with E. D correlates with E. However, A is independent of B. A and B are independent given C. B is independent of C. B and C are independent given A. C and E are independent given A, B, and D.
Hypothesis: There exists at least one collider (i.e., common effect) of A and B.
  • "Subtask 1"- The number of random variables and their names in the given premise in JSON format:
    Output: "number of random variables: 5, "names of random variables": ["A", "B", "C", "D", "E"]
  • "Subtask 2"- All the dependencies, conditional and unconditional independencies between all random variables extracted in "subtask 1":
    Output: {"All of Statistical Relations": {"Dependencies": [["A", "C"], ["A", "D"], ["A", "E"], ["B", "D"], ["B", "E"], ["C", "D"], ["C", "E"], ["D", "E"]], "Unconditional Independencies": [["A", "B"], ["B", "C"]], "Conditional Independencies": [["A", "B"], ["B", "C"], ["C", "E"]]}}
  • "Subtask 3"- The adjacency matrix of all random variables extracted in "subtask 1" where each random variable is treated as a node within a fully connected undirected graph:
    Output: {"A": {"A": 0, "B": 1, "C": 1, "D": 1, "E": 1}, "B": {"A": 1, "B": 0, "C": 1, "D": 1, "E": 1}, "C": {"A": 1, "B": 1, "C": 0, "D": 1, "E": 1}, "D": {"A": 1, "B": 1, "C": 1, "D": 0, "E": 1}, "E": {"A": 1, "B": 1, "C": 1, "D": 1, "E": 0}}
  • "Subtask 4"- Update the adjacency matrix extracted in the output of "subtask 3" based on the specified unconditional independencies between random variables. Each pair of variables that are declared independent should have their corresponding value set to zero in the adjacency matrix. - For each pair of variables listed as unconditional independent in "subtask 2", we set their corresponding entries in the adjacency matrix to 0. - We do not change any other entries except those specified by the independence.
    Output: {"A": {"A": 0, "B": 0, "C": 1, "D": 1, "E": 1}, "B": {"A": 0, "B": 0, "C": 0, "D": 1, "E": 1}, "C": {"A": 1, "B": 0, "C": 0, "D": 1, "E": 1}, "D": {"A": 1, "B": 1, "C": 1, "D": 0, "E": 1}, "E": {"A": 1, "B": 1, "C": 1, "D": 1, "E": 0}}
  • "Subtask 5"- Update the adjacency matrix in the output of "Subtask 4" based on the specified conditional independencies between random variables extracted in "subtask 2". Each pair of variables that are declared conditional independent should have their corresponding value set to zero in the adjacency matrix. - For each pair of variables listed as conditionally independent given other variable(s), we set their corresponding entries in the adjacency matrix to 0. - We do not change any other entries except those specified by the conditional independence.
    Output: {"A": {"A": 0, "B": 0, "C": 1, "D": 1, "E": 1}, "B": {"A": 0, "B": 0, "C": 0, "D": 1, "E": 1}, "C": {"A": 1, "B": 0, "C": 0, "D": 1, "E": 0}, "D": {"A": 1, "B": 1, "C": 1, "D": 0, "E": 1}, "E": {"A": 1, "B": 1, "C": 0, "D": 1, "E": 0}}
  • "Subtask 6": In this subtask, for each key value in the "Adjacency Matrix" mentioned in the output of "Subtask 5", the task is to extract only pairs of columns with values "1" step by step according to the provided steps and put it in "Candidates" dictionary: follow these steps: Step 1: Identify all rows (key values) in the matrix where there are at least two or more columns with the value "1". For each identified row, find all pairs of different columns where the values are "1". Ensure to exclude rows that do not contain any pairs from the results. Make sure to check each row individually and include all valid pairs for every row. Step 2: Display these pairs in a simplified JSON format, where each row name is a key, and the pair values are a list of column names that are identified in step 1. If there are no such rows, show an empty JSON object. Step 3: Do not include any rows with one or zero columns with "1" values in the output. Example output format: Candidates: { "row1 name": [["1value column1 name", "1value column2 name"], ["1value column1 name", "1value column3 name"], ["1value column2 name", "1value column3 name"]], "row2 name": [["1value column5 name", "1value column6 name"]], … } The values in each row have to be pairs and the output cannot be as follows: { "row1 name": [["1value column1 name"], ["1value column1 name"], ["1value column2 name", "1value column3 name"]], "row2 name": [["1value column5 name"]], … } Please provide only the desired output formatted exactly as shown in the example without any further explanation.
    Output: {"A": [["C", "D"], ["C", "E"], ["D", "E"]], "B": [["D", "E"]], "C": [["A", "D"]], "D": [["A", "B"], ["A", "C"], ["A", "E"], ["B", "C"], ["B", "E"], ["C", "E"]], "E": [["A", "B"], ["A", "D"], ["B", "D"]]}
  • "Subtask 7"- In this subtask, given the output of "Subtask 6", "Candidates", and "Unconditional Independencies" in "Subtask 2", the task is to identify and extract all the pairs in the "Candidates" that are also present in the "Unconditional Independencies" step by step according to the provided steps. follow these steps: 1- For each pair in the "Candidates" list, check if it is present in the "Unconditional Independencies" list. 3-Only keep all the pairs from "Candidates" that are also present in "Unconditional Independencies". If a pair in "Candidates" is found in "Conditional Independencies", keep it. 4-Remove any pairs in "Candidates" that are not found in "Conditional Independencies". If a pair in "Candidates" is not found in "Unconditional Independencies", remove it. 5-Output the result as the modified "Candidates" dictionary without any additional text or explanation. Only the updated "Candidates" dictionary and nothing else.
    Output: {"D": [["A", "B"], ["B", "C"]], "E": [["A", "B"]]}
  • "Subtask 8"- Given the adjacency matrix in the output of "Subtask 5" and the "All Pairs" list in the output of "Subtask 7", for each key-value pair ("R") in "All Pairs", we modify the adjacency matrix as follows: -For each key "R" and pair ("C1", "C2") in the candidates, change the values in key "R" and pairs ("C1", "C2") to zero. -Ensure that only the specified modifications are made, and all other entries in the adjacency matrix remain unchanged.
    Output: {"A": {"A": 0, "B": 0, "C": 1, "D": 1, "E": 1}, "B": {"A": 0, "B": 0, "C": 0, "D": 1, "E": 1}, "C": {"A": 1, "B": 0, "C": 0, "D": 1, "E": 0}, "D": {"A": 0, "B": 0, "C": 0, "D": 0, "E": 1}, "E": {"A": 0, "B": 0, "C": 0, "D": 1, "E": 0}}
  • "Subtask 9"- In this subtask, given the "Adjacency Matrix" mentioned in the output of "Subtask 8", evaluate the correctness of the hypothesis step by step according to the provided rules. Rules: 1- If matrix["R"]["C"] = 1 and matrix["C"]["R"] = 1, then the causal direction between "R" and "C" is undetermined. 2- If matrix["R"]["C"] = 1 and matrix["C"]["R"] = 0, then "R" is a direct cause of "C", or "C" is a direct effect of "R". 3- If matrix["R"]["C"] = 0 and matrix["C"]["R"] = 1, then "C" is a direct cause of "R", or "R" is a direct effect of "C". 4- Test each variable in the matrix, for instance, "R", "R" is a collider if the matrix entries in ["R", "C1"] = 0, ["C1", "R"] = 1, ["R", "C2"] = 0, and ["C2", "R"] = 1 Perform it step by step and provide the final "Yes" or "No" answer.
    Output: Step 1: According to the rules, a node is a collider (common effect) if the matrix entries in ["R", "C1"] = 0, ["R","C2"] = 0, ["C1", "R"] = 1 and ["C2", "R"] = 1.
    Step 2: According to rule 4, "R" is a collider if the matrix entries in ["R", "D"] = 0, ["R", "E"] = 0, ["D", "R"] = 1 and ["E", "R"] = 1.
    Step 3: Checking each variable in the matrix
    Node C: ["C", "A"] = 1 and ["A", "C"] = 1, so C is not a common effect of A and B.
    Node D: ["D", "A"] = 0 and ["A", "D"] = 1, and ["D", "B"] = 0 and ["B", "D"] = 1, so D is a common effect of A and B.
    Node E: ["E","A"] = 0 and ["A","E"] = 1, and ["E","B"] = 0 and ["B", "E"] = 1, so E is a common effect of A and B.
    Step 4: So, the hypothesis that there exists at least one collider (i.e., common effect) of A and B is correct. The answer is "Yes".

### v4

@@
-C2P: Featuring Large Language Models with Causal Reasoning
+
 Abstract
-Causal reasoning is the primary bottleneck that Large Language Models (LLMs) must overcome to attain human-level intelligence.
-To address this, we introduce the Causal Chain of Prompting (C2P) as the first reasoning framework that equips current LLMs with causal reasoning capabilities.
-C2P operates autonomously, avoiding reliance on external tools or modules during both the causal learning and reasoning phases, and can be seamlessly implemented during the training or fine-tuning of LLMs.
-Experimental results across various benchmark datasets demonstrate a significant improvement in causal learning and subsequent reasoning accuracy of LLMs.
-We illustrate how C2P enhances LLMs’ ability to causally reason in real-world scenarios, addressing complex problems in fields such as healthcare, medicine, economics, education, social sciences, environmental science, and marketing.
-With few-shot learning, GPT-4 Turbo using C2P with as few as six examples achieves significant performance improvements, boasting over a $33\%$ increase in reasoning accuracy over the most state-of-the-art LLMs, which perform nearly randomly in similar circumstances.
-This demonstrates the transformative potential of integrating C2P into LLM training or fine-tuning processes, thereby empowering these models with advanced causal reasoning capabilities.
+Causal reasoning is one of the primary bottlenecks that Large Language Models (LLMs) must overcome to attain human-level intelligence.
+Recent studies indicate that LLMs display near-random performance on reasoning tasks.
+To address this, we introduce the Causal Chain of Prompting ($\text{C}^{2}\text{P}$), a reasoning framework that aims to equip current LLMs with causal reasoning capabilities as the first framework of its kind operating autonomously without relying on external tools or modules during both the causal learning and reasoning phases.
+To evaluate the performance of $\text{C}^{2}\text{P}$, we first demonstrate that reasoning accuracy improved by over $30.7\%$ and $25.9\%$ for GPT-4 Turbo and LLaMA 3.1, respectively, when using our framework, compared to the same models without $\text{C}^{2}\text{P}$ on a synthetic benchmark dataset.
+Then, using few-shot learning of the same LLMs with $\text{C}^{2}\text{P}$, the reasoning accuracy increased by more than $20.05\%$ and $20.89\%$, respectively, with as few as ten examples, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on the same dataset.
+To evaluate $\text{C}^{2}\text{P}$ in realistic scenarios, we utilized another benchmark dataset containing natural stories across various fields, including healthcare, medicine, economics, education, social sciences, environmental science, and marketing.
+The results show improved reasoning when $\text{C}^{2}\text{P}$ is applied, compared to cases where our framework is not used, which often leads to random and hallucinated responses.
+By showing the improved performance of few-shot learned GPT-4 Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$, we demonstrate the generalizability of our framework.
 1 Introduction
-Recent advancements in Large Language Models (LLMs) have impacted existing AI paradigms and heightened expectations regarding AI’s capabilities (Achiam et al., 2023; Brown et al., 2020).
-Figure 1: Example of the standard prompting vs few-shot learned GPT-4 with C2P in open problem in astrophysics (Pasquato et al., 2023)
-Despite significant architectural differences in LLMs, they generally produce outputs based on the most likely results learned from vast amounts of training data (Vaswani et al., 2017).
-This enables them to acquire extensive knowledge ranging from common sense to specialized domains such as mathematics and science (Jiralerspong et al., 2024).
-Despite this, the inefficiency of LLMs in addressing causal reasoning questions remains their primary bottleneck and simple tasks can completely break down reasoning in state-of-the-art LLMs (Nezhurina et al., 2024).
-Additionally, studies such as Kalai & Vempala (2023); Xu et al.
-(2024) have demonstrated that, despite the training data containing numerous examples of interventions, outcomes, and explanations, as well as similar tasks, hallucinatory responses persist and there is a lack of causal reasoning capability.
-As a result, while they may talk causality, they are not causal (Zečević et al., 2023).
-This deficiency represents a fundamental drawback of LLMs as AI systems compared to human intelligence, which goes beyond mere correlations and depends on causal relationships for decision-making (Penn & Povinelli, 2007; Anwar et al., 2024).
-Recently, answering cause and effect questions with LLMs has gained extensive interest (Shin et al., 2020; Jin et al., 2023b; Ashwani et al., 2024).
-To address this issue, it is important to note that LLMs have been utilized in conjunction with external tools to extract causal structures, as demonstrated in (Jiralerspong et al., 2024).
-However, their architectures lack specialized modules specifically designed to enhance the understanding of cause-and-effect relationships within their outputs (Wang et al., 2023; Imani et al., 2023).
-Aside from studies that reason causally based on the knowledge that is already in their training data (Petroni et al., 2019; Jiang et al., 2020) and the ones that use LLMs in causality (Kıcıman et al., 2023; Zhang et al., 2023; Feder et al., 2024; Khatibi et al., 2024), chain-of-thought prompting is presented in (Wei et al., 2022) as one of the initial attempts in enhancing the reasoning in LLMs that shows improvement based on the data of the given query.
-However, LLMs still struggle with rigorous numerical and abstract reasoning among many other tasks (Xu et al., 2023).
-For example, a recent work, Causal Reasoning Assessment Benchmark (CRAB, Romanou et al., 2023), is designed to evaluate the causal understanding of events in real-world narratives.
-This study demonstrated that most systems perform poorly in identifying cause-and-effect tasks.
-Similarly, in (Jin et al., 2023b), the CORR2CAUSE dataset is introduced and demonstrated that current models often perform no better than random chance when tasked with causal questions.
-Following that, in (Jin et al., 2023a), the CLADDER dataset is introduced to asses Average Treatment Effects with LLMs and it is demonstrated that these models struggle with causal tasks.
-In their study, with the implementation of the proposed framework, CAUSALCoT, progress has been made in evaluating the average treatment effect where LLMs are provided with a collection of causal graphs and various types of queries (associational, interventional, and counterfactual), such as those included in the CLADDER dataset.
+Recent advancements in Large Language Models (LLMs) have impacted existing AI paradigms and raised expectations regarding AI’s capabilities (Achiam et al., 2023; Brown et al., 2020).
+Figure 1: The 5 successive subtasks of $\text{C}^{2}\text{P}$ applied to Black Holes and their Host Galexis example presented in (Pasquato et al., 2023).
+Central Density(CD), Black Hole Mass(BHM), Velocity Density(VD), Bulge Stellar Mass(BSM), Effective Radius(ER)
+LLMs generally produce outputs based on the most likely results learned from vast amounts of training data (Vaswani et al., 2017).
+This enables them to acquire extensive knowledge, ranging from common sense to specialized domains such as mathematics and science (Jiralerspong et al., 2024).
+Numerous examples of interventions, outcomes, and explanations are included in their training of LLMs to reduce hallucinations and improve reasoning.
+However, despite significant architectural differences, hallucinatory responses still occur and true causal reasoning remains lacking (Kalai & Vempala, 2023; Xu et al., 2024).
+As a result, although models may appear to reason causally, they do not engage in a genuine causal reasoning process (Zečević et al., 2023).
+This deficiency represents a fundamental limitation of LLMs as AI systems compared to human intelligence, which is based on causal reasoning rather than simple associations for decision making (Penn & Povinelli, 2007; Anwar et al., 2024).
+Judea Pearl introduced "The Ladder of Causation" in (Pearl & Mackenzie, 2018) that easily addresses the reason for this deficiency.
+This ladder includes three main levels.
+At the first level, association, only patterns and dependencies are identified, and why things are related cannot be answered.
+The second level, intervention, involves understanding cause-and-effect by predicting outcomes of actions or changes, asking "What happens if I do this?" The highest level, counterfactuals, deals with imagining alternative scenarios, asking "What would have happened if…?".
+This framework illustrates how reasoning evolves from pattern recognition to understanding and predicting causal relationships, highlighting the differences at each level.
+Recently, studies on the causal reasoning capabilities of LLMs have garnered significant interest.
+Most of these focus on evaluating the reasoning abilities of LLMs, while a smaller number aim to enhance this feature.
+Additionally, some studies explore the application of LLMs in the field of causality.(Kıcıman et al., 2023; Zhang et al., 2023; Feder et al., 2024; Khatibi et al., 2024).
+The inefficiency of LLMs in reasoning has been extensively studied from various perspectives, as demonstrated in works such as (Xu et al., 2023; Romanou et al., 2023; Jin et al., 2023a; Liu et al., 2024a; Hobbhahn et al., 2022).
+Similarly, Jin et al.
+(2023b) introduces the CORR2CAUSE dataset, revealing that current models often perform no better than random chance when answering causal questions.
+Most recently, simple tasks have been shown to completely break down the reasoning abilities of state-of-the-art LLMs (Nezhurina et al., 2024).
+Additionally, studies such as (Petroni et al., 2019; Jiang et al., 2020) aimed to reason causally based on the knowledge already present in the training data of LLMs, which is why LLMs are referred to as "causal parrots" in (Zečević et al., 2023).
+The reason for these behaviors is discussed in (Wang et al., 2023; Imani et al., 2023).
+As one of the initial attempts to enhance reasoning in LLMs, chain-of-thought prompting is presented in (Wei et al., 2022b), showing improvement based on the data from the given query.
+Despite CoT’s potential to enhance reasoning performance, the analysis in Bao et al.
+(2024) reveals a pattern: correct answers often follow incorrect CoTs, and vice versa.
+As another approach, LLMs have been utilized in conjunction with external tools to extract causal structures, as demonstrated in (Jiralerspong et al., 2024).
 More recently, in (Ashwani et al., 2024), a novel architecture called the Context-Aware Reasoning Enhancement with Counterfactual Analysis (CARE-CA) framework is presented to enhance causal reasoning and explainability.
@@
 Several other works at the intersection of causal inference and LLMs are discussed in an extensive survey by Liu et al.
-(2024).
-As argued by Pearl (1995), causal Directed Acyclic Graphs (DAGs), along with d-separation, allow for the investigation of cause-and-effect relationships without relying on structural equation models in computational studies.
-Inspired by Pearl’s foundational work, we propose a novel framework, named Causal Chain of Prompting (C2P), to address the inefficiencies of LLMs in handling causal queries (see Figure 1 as an example).
-We show that by identifying the adjacency matrix, equivalent to and instead of the causal DAG in Pearl’s work, the causal relationships among the variables in the premise’s cause-and-effect relationships can be effectively reasoned within the context of language models.
-Contrary to the existing research in addressing weaknesses of LLMs in causal reasoning, the C2P operates autonomously, avoiding reliance on external tools or modules during both the learning and reasoning phases to answer causal questions.
-Additionally, C2P can easily be implemented in the training or fine-tuning process of LLMs.
-C2P includes five simple main sub-tasks, as follows: (1) Prompting for extracting the random variables from the provided data.
-(2) Prompting for extracting all conditional and unconditional relations and cause and effect relations specifically mentioned among the random variables.
-(3) Prompting to create the initial adjacency matrix with values 1 for all elements except the diagonal elements and the ones corresponding to effect-cause relations (the cause and effects elements are also one).
-(4) Prompting of the conditional and unconditional independencies and identification of the colliders, step by step, to extract the causal adjacency matrix; and, (5) Prompting for reasoning questions or hypotheses.
-To evaluate the accuracy and reliability of implementing the C2P on LLMs, we initially assess it using publicly available synthetic datasets such as (Jin et al., 2023b).
-Subsequently, we evaluate it in more realistic and complex scenarios in real-world problems presented in (Pearl & Mackenzie, 2018) and (Pasquato et al., 2023).
-Moreover, we present results on few-shot learning with C2P on both synthetics and realistic scenarios.
+(2024b).
+The main drawbacks of the existing frameworks aiming to equip LLMs with reasoning are their reliance on external modules, the need for extensive information to function, and their very low accuracy.
+In this paper, we propose the Causal Chain of Prompting ($\text{C}^{2}\text{P}$), a reasoning framework for LLMs designed to enhance reasoning skills by climbing the causality ladder and answering reasoning questions.
+Unlike existing methods, $\text{C}^{2}\text{P}$ operates autonomously, without relying on external tools or modules during the learning and reasoning phases.
+It can be easily implemented in the few-shot to improve reasoning in causal questions.
+Importantly, $\text{C}^{2}\text{P}$ extracts the causal relation based on associations mentioned in a given premise.
+$\text{C}^{2}\text{P}$ is inspired by Pearl’s foundational work, which argues that causal Directed Acyclic Graphs (DAGs), along with d-separation, enable the investigation of cause-and-effect relationships without relying on structural equation models in computational studies (Pearl, 1995).
+Based on this, we demonstrate that by identifying the adjacency matrix of causal relationships among variables in the premise—similar to, but distinct from, the causal DAG in Pearl’s framework—a reasoning query can be effectively answered.
+This framework includes five simple main subtasks, as follows: (1) Prompting to extract random variables from the provided data.
+(2) Prompting to extract all conditional and unconditional associations, as well as cause-and-effect relations specifically mentioned among the random variables.
+(3) Prompting to create the initial adjacency matrix with values of 1 for all elements except the diagonal elements and those corresponding to effect-cause relations (the cause-and-effect elements are also set to 1).
+(4) Prompting conditional and unconditional independencies and identification of colliders, step by step, to extract the causal adjacency matrix.
+(5) Prompting for reasoning questions or hypotheses.
+(see Fig.
+1).
+To evaluate the accuracy and reliability of implementing the $\text{C}^{2}\text{P}$ on LLMs, we initially assess it using publicly available benchmark synthetic and "Natural Story" datasets, such as those in (Jin et al., 2023b).
+To demonstrate the practical applicability of $\text{C}^{2}\text{P}$, we present the results of few-shot learned LLMs with $\text{C}^{2}\text{P}$ in both synthetic and real-world scenarios.
+Subsequently, we evaluate it in more realistic and complex scenarios found in real-world problems presented in (Pasquato et al., 2023).
 Contributions.
-In this work, we present several important contributions to facilitate causal reasoning in language models. Concretely,
+In this work, we present several contributions to facilitate causal reasoning in large language models. Concretely,
   1. 1.
-    We introduce the C2P framework as the first reasoning framework to equip LLMs with the ability of causal reasoning within real-world scenarios, without relying on external tools.
+    We introduce the $\text{C}^{2}\text{P}$ framework as the first reasoning framework to equip LLMs with causal reasoning within real-world scenarios, without relying on external tools.
   2. 2.
-    Through extensive experiments with our framework, we demonstrate a significant improvement of LLMs in causal reasoning in various benchmarks.
-Additionally, we examine the performance of the C2P framework on more complex and real-world scenarios in various domains.
+    Through extensive experiments with our framework, we demonstrate significant improvements in the causal reasoning abilities of LLMs across various benchmark datasets, as well as in more complex, real-world scenarios across multiple domains.
   3. 3.
-    We perform few-shot learning experiments in section 4 with GPT-4 Turbo using our C2P framework and show how integrating C2P during the training or fine-tuning of LLMs can revolutionize existing models.
-This approach equips them with causal reasoning capabilities akin to the transformative impact of ‘Chain of Tought’ (Wei et al., 2022) on LLMs, as discussed by Chung et al.
-(2024).
-The codes for sections 3 and 4 are publicly available at https://anonymous.4open.science/r/C2P-5C2A/.
+    We demonstrate the generality of $\text{C}^{2}\text{P}$ by implementing few-shot learning for both GPT-4 Turbo and LlaMA 3.1, obtaining significant improvements in reasoning.
+The code is publicly available at https://github.com/abmbagheri/c2p-Featuring-Large-Language-Models-with-Causal-Reasoning.git.
 2 Preliminaries on Causal Learning and Reasoning
 To lay the groundwork for our framework, understanding the fundamental concept of cause and effect is paramount.
-A foundational and widely acknowledged principle in causality is the one proposed by Reichenbach, which posits the following:
-Principle 1 (Common cause principle (Reichenbach, 1991))
-[…]
-Markov Property: The Markov property in a DAG $G$ states that each node $X_{i}$ is conditionally independent of its non-descendants given its parents, namely $X_{i}\perp\!\!\!\perp\text{NonDe}(X_{i})|\text{Pa}(X_{i})$, where $\text{NonDe}(X_{i})$ denotes the non-descendants of $X_{i}$ excluding itself, and $\text{Pa}(X_{i})$ denotes the parents of $X_{i}$.
-Using the Markov property, we can factorize the joint distribution of all the nodes in the graph into $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i}|\text{Pa}(X_{i}))$.
-To infer the causal graph from probability distributions, a common assumption is faithfulness, namely the validity to infer all the d-separation sets in the graph from the independence relations in the probability distribution.
-In our work, we also take this broadly taken assumption which holds for most real-world scenarios.
-Directed Acyclic graph: A graph $G$ is called a partially directed acyclic graph (PDAG) if there is no directed cycle, that is, if there is no pair $(j,k)$ with directed paths from $j$ to $k$ and from k to $j$.
+A widely acknowledged principle in causality is the principle proposed by Reichenbach, which posits the following.
+Common cause principle (Reichenbach, 1991) If two random variables $X_{1}$ and $X_{2}$ are statistically dependent, i.e., $X_{1}\not\!\perp\!\!\!\perp X_{2}$, then there exists a third variable $X_{3}$ that causally influences both.
+(As a special case, $X_{3}$ may coincide with either $X_{1}$ or $X_{2}$.) Furthermore, this variable $X_{3}$ screens $X_{1}$ and $X_{2}$ from each other in the sense that given $X_{3}$, they become independent, $X_{1}\perp\!\!\!\perp X_{2}|X_{3}$.
+do calculus (Pearl, 1995): do-calculus, developed by Judea Pearl et al., is a set of rules used to transform and manipulate causal expressions within causal diagrams (or graphical models).
+do-calculus is a formal tool used to reason in causal relationships from a mixture of experimental and observational data.
+do-calculus consists of three main rules that allow one to rewrite expressions involving interventions (typically represented as $do(x)$, indicating an intervention to set variable $X$ to value $x$).
+These rules are crucial for determining the identifiability of causal effects from data, allowing researchers to reason about causal relationships using a combination of experimental and observational data.
+The rules are provided in Appendix A.1.
+d-separation (Pearl, 1995).
+d-Separation is a criterion used in Bayesian network analysis to determine whether a set of variables $X_{1}$ is independent of another set of variables $X_{2}$, given a third set of variables $X_{3}$.
+This concept is foundational in understanding the flow of causal effects in graphical models and helps in deciding whether a path between two variables is "blocked" or not by conditioning on other variables.
+According to the d-separation, a path between two variables is blocked if it includes an intermediate variable that is a collider and is not conditioned on or a non-collider that is conditioned on.
+Here, a collider is a variable that has arrows inward from two other nodes (i.e., $X_{1}\rightarrow X_{2}\leftarrow X_{3}$), whereas a non-collider does not meet this criterion.
+Directed Acyclic Graphs.
+A graph $G$ is called a Partially Directed Acyclic Graph (PDAG) if there is no directed cycle, that is, if there is no pair $(X_{j},X_{k})$ with directed paths from $X_{j}$ to $X_{k}$ and from k to $j$.
 $G$ is called a Directed Acyclic Graph (DAG) if it is a PDAG and all edges are directed.
-Markov Equivalence of Graphs: We denote two DAGs as Markov equivalent if they induce the same joint distribution $P(\mathbf{X})$.
-The set of DAGs that are Markov equivalent to each other is called a Markov equivalence class (MEC).
-Causal graphs in the same MEC can be easily identified since they have the same skeleton (i.e., undirected edges) and V-structures (i.e., structures in the form of $X_{1}\rightarrow X_{2}\leftarrow X_{3}$ where $X_{1}$ and $X_{3}$ are not connected).
-The be able to perform any level of causal reasoning, the Ladder of Causation is proposed with three main rungs, “seeing”, “doing”, and “imagining”, in (Pearl & Mackenzie, 2018).
-To fulfill all rungs of the ladder, more information on the causal structure is needed, which can be mostly provided with structural causal models (SCM) (Assuming all the assumptions are satisfied) (Bareinboim et al., 2022).
-The current causal discovery methods are divided into two groups, constraint-based algorithm such as the PC algorithm, which has a PDAG as an output, that represents the MEC of the true underlying graph and is the best result such methods can extract.
-Second is score-based methods such as NOTEARS (Zheng et al., 2018), GOLEM (Ng et al., 2020), and DAGMA (Bello et al., 2022), among many others, that extract a DAG that mostly fits the data and demonstrated very high accuracy in extracting Bayesian networks.
-Score-based methods require solving an optimization problem based on the numerical values of variables, rendering them impractical for implementation in LLM models.
-PC algorithm: Following these materials, the PC algorithm is developed based on Reichenbach’s common cause principle and Markov property as follows for a causal mechanism with $p$ random variables (Spirtes et al., 2001);
+Markov Property.
+The Markov property in a DAG $G$ states that each node $X_{i}$ is conditionally independent of its non-descendants, given its parents.
+In other words, $X_{i}\perp\!\!\!\perp\text{NonDe}(X_{i})|\text{Pa}(X_{i})$, where $\text{NonDe}(X_{i})$ represents the non-descendants of $X_{i}$, excluding itself, and $\text{Pa}(X_{i})$ represents the parents of $X_{i}$.
+It helps factorize the distribution of all graph nodes as $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i}|\text{Pa}(X_{i}))$.
+Faithfulness.
+This assumption ensures that all the d-separation sets in the graph can be inferred from the independence relations in the distribution.
+In the sequel, we assume faithfulness, a widely used assumption in causal discovery (Spirtes et al., 2001).
+Markov Equivalence of Graphs.
+Two DAGs are Markov equivalent if they generate the same joint distribution, $P(\mathbf{X})$.
+A set of DAGs that are Markov equivalent is referred to as a Markov equivalence class (MEC).
+Causal graphs within the same MEC are easily recognizable because they share the same skeleton (i.e., the same undirected edges) and the same V-structures (i.e., configurations like $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, where $X_{1}$ and $X_{3}$ are not directly connected).
+Ladder of Causation (Pearl & Mackenzie, 2018).
+To perform any level of causal reasoning, the Ladder of Causation proposes three main levels: “seeing or association,” “doing or intervention,” and “imagining or counterfactual”.
+The questions that can be answered in Association (Seeing) are mostly similar to "What is happening?" This is the most basic level where we observe patterns and correlations between variables.
+At this level, we can only say that two things are related or tend to occur together, but we cannot explain why.
+Machine learning models and statistical methods that rely on pattern recognition (like LLMs) often operate at this level.
+In the second level, Intervention (Doing), questions such as "What happens if I do something?" can be addressed.
+At this level, we can go beyond mere association and ask about the effect of an action or intervention.
+This requires understanding the cause-and-effect relationships.
+To reach the third level of the ladder, more information is needed on the causal structure, which can mainly be provided with structural causal models (SCMs), assuming that all assumptions are satisfied (Bareinboim et al., 2022).
+However, studies such as (Spirtes et al., 2001) have shown that Level 2 can be reached (up to equivalence classes) using only observational data.
+Note that the current causal discovery methods are primarily divided into two groups: (i) Constraint-based algorithms such as the PC algorithm (Spirtes et al., 2001), which has a PDAG as an output, which represents the MEC of the true underlying graph and is the best outcome that these methods can achieve; (ii) Score-based methods such as GES (Chickering, 2002), NOTEARS (Zheng et al., 2018), GOLEM (Ng et al., 2020), DAGMA (Bello et al., 2022), and TOPO (Deng et al., 2023), among many others, which extract a DAG that mostly fits the data and demonstrated high accuracy in extracting Bayesian networks.
+Nevertheless, score-based methods require the solution of a numerical optimization problem, making them difficult to integrate with LLMs.
+The PC algorithm (Spirtes et al., 2001). The PC algorithm is developed based on Reichenbach’s common cause principle and the Markov property, its steps can be described as follows:
   1. (i)
@@
   4. (iv)
-    For each pair of variables $X_{1},X_{2}$ having an edge between them, and for each pair of variables $\{X_{3},X_{4}\}$ with edges both connected to $X_{1}$ or both connected to $B$, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|\{X_{3},X_{4}\}$.
+    For each pair of variables $X_{1},X_{2}$ having an edge between them, and for each pair of variables $\{X_{3},X_{4}\}$ with edges both connected to $X_{1}$ or both connected to $X_{2}$, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|\{X_{3},X_{4}\}$.
   5. (v)
-    For each triple of variables $(X_{1},X_{2},X_{3})$ such that $X_{1}$ and $X_{2}$ are adjacent, $X_{2}$ and $X_{3}$ are adjacent, and $A$ and $C$ are not adjacent, orient the edges $X_{1}-X_{2}-X_{3}$ as $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, if $X_{2}$ was not in the set conditioning on which $X_{1}$ and $X_{3}$ became independent and the edge between them was accordingly eliminated.
+    For each triple of variables $(X_{1},X_{2},X_{3})$ such that $X_{1}$ and $X_{2}$ are adjacent, $X_{2}$ and $X_{3}$ are adjacent, and $X_{1}$ and $X_{3}$ are not adjacent, orient the edges $X_{1}-X_{2}-X_{3}$ as $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, if $X_{2}$ was not in the set conditioning on which $X_{1}$ and $X_{3}$ became independent and the edge between them was accordingly eliminated.
 We call such a triple of variables a V-structure.
@@
 This is called orientation propagation.
-d-separation and do calculus (Pearl, 1995): do-calculus, developed by Judea Pearl and others, is a set of rules used to transform and manipulate causal expressions within causal diagrams (or graphical models).
-It is a generalization of do-calculus, which is itself a formal tool used to reason in causal relationships from a mixture of experimental and observational data.
-do-calculus consists of three main rules that enable one to rewrite expressions involving interventions (typically represented as $do(x)$, indicating an intervention to set variable $X$ to value $x$):
-  1. 1.
-    Insertion/deletion of observations: If a variable $X_{2}$ is independent of $X_{1}$ given a set of other variables $X_{3}$, then the conditional distribution of $X_{2}$ given $do(X_{1})$ and $X_{3}$ is the same as the conditional distribution of $X_{2}$ given $X_{3}$ alone.
-This rule can be applied when $X_{1}$ and $X_{2}$ are d-separated by $X_{3}$ in the graph modified by removing all arrows pointing into $X_{1}$.
-  2. 2.
-    Action/observation exchange: If $X_{2}$ is independent of $X_{1}$ given $X_{3}$ and another set $W$ where $X_{1}$ is intervened, then the graph can be modified to reflect this exchange.
-  3. 3.
-    Insertion/deletion of actions: This rule states that if intervening on $X_{1}$ does not affect $X_{2}$ after accounting for $X_{3}$, the intervention can be ignored in the expression.
-These rules are crucial for determining the identifiability of causal effects from data, allowing researchers to reason in causal relationships from a combination of experimental and observational data.
-d-Separation is a criterion used in the analysis of Bayesian networks to determine whether a set of variables $X_{1}$ is independent of another set of variables $X_{2}$, given a third set of variables $X_{3}$.
-This concept is foundational in understanding the flow of causal influence in graphical models and helps in deciding whether a path between two variables is "blocked" or not by conditioning on other variables.
-The basic rule of d-separation is: A path between two variables is blocked if it includes an intermediate variable that is a collider and is not conditioned on or a non-collider that is conditioned on.
-Here, a collider is a variable that has inward arrows from two other nodes (i.e., $X_{1}\rightarrow X_{2}\leftarrow X_{3}$), whereas a non-collider does not meet this criterion.
-3 Developing Causal Chain of Prompting
-To develop C2P framework, we aim to extract the adjacency matrix of variables in a given premise.
-The adjacency matrix acts as equivalent to and instead of the causal DAG in Pearl’s work leading to answering the causal reasoning question.
+3 The Causal Chain of Prompting
+To develop our framework, our aim is to extract the adjacency matrix of variables on a given premise.
+The adjacency matrix acts as equivalent to and instead of the causal DAG in Pearl’s work, leading to answering the causal reasoning question.
 In developing the Causal Chain of Prompting, we use a few steps similar to the PC algorithm for the given premise.
-3.1 Causal Chain of Prompting (C2P)
-The C2P frameworks consist of 5 main sub-tasks as follows, for learning and reasoning cause and effect relation for the given premise;
-  1. 1.
-    Prompting to extract the random variables in the provided data.
-  2. 2.
-    Prompting to extract all the cause and effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise.
-  3. 3.
-    Prompting to create an initial adjacency matrix where all elements are 1, except for the diagonal elements and the elements corresponding to the cause-and-effect relationships specifically mentioned in the given premise (extracted in sub-task 2).
-  4. 4.
-    Prompting of the conditional and unconditional independency valuation and identifying the colliders, step by step, to extract the causal PDAG.
-  5. 5.
-    Prompting for cause-and-effect questions or hypotheses.
-Each sub-task in C2P includes multiple steps and overall, to execute the frameworks, 9 main prompts have to be completed. The exact prompts for all the steps are provided in the Appendix, A.1.
-3.2 Few-shot learning agenda
-LLMs have powerful zero-shot capabilities, yet they might struggle with complex tasks with insufficient examples in the training process.
-In these situations, few-shot learning is a versatile and efficient technique for in-context learning that can be used to quickly adapt LLMs to new tasks and significantly enhance their performance (Min et al., 2022; Touvron et al., 2023).
-This approach involves providing several examples with desired answers, which condition the LLMs to produce the correct response for new instances with similar patterns.
-The few-shot learning process of C2P is similar to the previous subsection, with the prompts being based on the more abstract prompts presented there (Due to the token limitation).
-The prompts and an example of the given story are included in the Appendix, A.2.
-Depending on the token limitation of employed LLMs, the number of given examples (shots) varies.
-Comparing the token limitation of GPTs, we use GPT-4 Turbo, which has a 30000 token limit.
-Steps from other constraint-based methods can be used instead or along with some steps of the PC algorithm in C2P methods, comprehensively discussed in (Glymour et al., 2019).
-For instance, steps in the FCI method can be used in cases where the causal sufficiency assumption is violated, i.e., latent variables and selection bias may be present (Spirtes et al., 1995).
-Additionally, it is important that the classifiers proposed by Ceraolo et al.
-(2024) that formalize the definition of causal questions and establish a taxonomy for finer-grained classification can be used before our framework to identify the causal question as a prerequisite of our method.
-4 Results of the experiment for featuring LLMs with Causal Reasoning
+3.1 Causal Chain of Prompting ($\text{C}^{2}\text{P}$)
+The $\text{C}^{2}\text{P}$ framework consists of five main subtasks as follows, for learning and reasoning about cause and effect relations for the given premise:
+  • Subtask 1: Prompting to extract the random variables in the provided data.
+  • Subtask 2: Prompting to extract all the cause-and-effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise.
+  • Subtask 3: Prompting to create an initial adjacency matrix where all elements are 1, except for the diagonal elements and those corresponding to the cause-and-effect relationships specifically mentioned in the given premise (extracted in subtask 2).
+  • Subtask 4: Prompting of conditional and unconditional independence evaluation and identification of colliders to extract the causal PDAG.
+  • Subtask 5: Prompting for cause-and-effect questions or hypotheses.
+Each subtask in $\text{C}^{2}\text{P}$ can include one or multiple steps (prompts).
+In general, to execute the framework, 9 main steps must be completed.
+Subtask 1 is completed in Step 1.
+Subtask 2 is accomplished with Step 2.
+Subtask 3 is achieved with Step 3.
+To perform Subtask 4, 5 steps must be applied.
+Step 4 first eliminates all unconditional independencies achieved in Subtask 2 and Step 5 then removes all conditional independencies extracted in Subtask 2.
+Step 6 identifies potential nodes that can act as colliders.
+Step 7 confirms whether the nodes identified in Step 6 are colliders.
+Step 8 updates the adjacency matrix, resulting in the final adjacency matrix of a causal structure.
+To perform Subtask 5, the causal question is asked in Step 9.
+The exact prompts for all the steps are provided in Appendix A.3 and code repository.
+3.2 Few-shot learning
+LLMs have powerful zero-shot capabilities, yet they struggle with complex tasks because of their engineering designs and insufficient examples in the training process.
+In such situations, few-shot learning is a versatile and efficient technique for in-context learning, which can be used to quickly adapt LLMs to new tasks and significantly enhance their performance (Min et al., 2022; Touvron et al., 2023).
+This approach involves providing several examples with desired answers to condition the LLMs to produce correct responses for new instances with similar patterns.
+The few-shot learning process of $\text{C}^{2}\text{P}$, as described in the previous subsection, is based on more abstract prompts (due to token limitations).
+The prompts and an example of the given story are included in Appendix A.4.
+Depending on the token limitations of the employed LLMs, the number of examples (shots) may vary.
+For GPT-4 Turbo, which has a 30,000 token limit, we were able to include ten examples.
+For LLaMA 3.1, we used the same examples in the few-shot learning process as well.
+It is important to note that steps from other constraint-based methods can either replace or complement the steps of the PC algorithm in $\text{C}^{2}\text{P}$ methods, as comprehensively discussed in (Glymour et al., 2019).
+For instance, the steps in the FCI method can be used in cases where the causal sufficiency assumption is violated, i.e., where latent variables and selection bias may be present (Spirtes et al., 1995).
+Additionally, the classifiers proposed by Ceraolo et al.
+(2024) formalize the definition of causal questions and establish a taxonomy for finer-grained classification.
+These classifiers can be used before our framework to identify the causal question as a prerequisite of our method.
+4 Experiments
+4.1 Datasets
+The experiments are divided into two data groups: a synthetic dataset and a real-world scenarios dataset.
+The CORR2CAUSE dataset:
+Introduced in (Jin et al., 2023b), the CORR2CAUSE dataset serves as a benchmark to assess the ability of LLMs to respond to reasoning queries.
+Tab.
+LABEL:table01 shows the details of the samples in this dataset.
+The process of creating CORR2CAUSE is as follows: First, select the number $N$ of variables (Step 1) and generate all unique DAGs with $N$ nodes (Step 2).
+Next, gather all d-separation sets from these graphs to identify MECs (Step 3).
+In Step 4, formalize the data by associating each MEC with its corresponding causal graphs.
+For each MEC, construct a correlation statement based on the statistical relations within the MEC, hypothesize a causal relationship between two variables, and assign a validity "Yes" if the hypothesis holds for all causal graphs in the MEC, or "No" if the hypothesis does not necessarily apply to all graphs in the MEC.
+Finally, introduce the verbalization process.
+An example of a premise and its corresponding hypothesis in the CORR2CAUSE dataset is as follows:
+Premise: Suppose that there is a closed system of 3 variables, A, B and C.
+All statistical relations among these 3 variables are as follows: A correlates with C.
+B correlates with C.
+However, A is independent of B.
+Hypothesis: A directly affects C.
+Table 1: Information on CORR2CAUSE samples, different variable numbers, and responses.
+Scenario Varible 3 varibles 4 varibles 5 varibles 6 varibles Sum
+CORR2CAUSE test data
+All samples 45 72 518 522 1157
+Samples with "No" answers 30 57 449 402 938
+Samples with "Yes" answers 15 15 69 120 338
+CORR2CAUSE train data
+All samples 0 576 7524 197634 205734
+Samples with "No" answers 0 529 6542 160615 167686
+Samples with "Yes" answers 0 47 982 37019 38048
+The Natural Stories dataset:
+The Natural Stories dataset is also introduced in (Jin et al., 2023b) to assess the reasoning capabilities of LLMs in realistic scenarios.
+This dataset builds upon the CORR2CAUSE dataset as a foundation for future extensions in various contexts, such as instantiating variables with real-world phenomena and placing the narratives in more natural settings.
+For instance, the rule "correlation does not imply causation" can be illustrated using ice cream sales and swimming pool attendance as variables, by arguing that ice cream sales do not necessarily impact swimming pool attendance, as both could be influenced by a third factor, such as hot weather.
+The Natural Stories data presented in (Jin et al., 2023b) is not open-source.
+However, generating such data is straightforward.
+By providing an example from the CORR2CAUSE dataset along with the corresponding Natural Stories created by a human or existing example on the web, GPT-4 can generate realistic, everyday narratives.
+For more detailed examples of generating such stories, refer to the code directory accompanying this paper.
+A natural story generated by GPT-4, inspired by the previous example, could be:
+Premise: Let’s consider three factors: eating junk food, obesity, and watching television.
+There is a correlation between eating junk food and obesity, and between watching television and obesity.
+However, eating junk food and watching television are independent from each other.
+Hypothesis: Eating junk food directly affects obesity.
+More details and examples are provided in the code repository of this study.
+An example on supermassive black holes.
+The data and results on the coevolution of supermassive black holes (SMBHs) and their host galaxies, presented in (Pasquato et al., 2023), are used as a real-world example in its verbalized form.
+The verbalized information is as follows:
+Premise: In the (co)evolution of supermassive black holes (SMBHs) and their host galaxies, the data supports the following observations.
+Existing studies show that with a change in central density, there is no significant change in bulge stellar mass.
+However, there is a decrease in black hole mass with higher central density.
+Additionally, when black hole mass is fixed, central density and velocity dispersion change simultaneously.
+Conversely, higher velocity dispersion or effective radius results in lower bulge stellar mass, while velocity dispersion and effective radius do not change simultaneously.
+Hypothesis: Does central density affect black hole mass?
 4.2 Experimental setup
-The experimental setup is divided into two data groups: synthetic dataset and realistic scenarios.
-In a first synthetic dataset, and similar to the experimental setup of (Jin et al., 2023b), we set up the following list of LLMs for the experiments on the CORR2CAUSE dataset.
-Various types of LLMs are tested in (Jin et al., 2023b) and we only include powerful LLMs in our experiment to avoid reproducing the same results.
-To test existing LLMs, we first include two BERT-based NLI models in the transformers library (Wolf et al., 2020): BART (Lewis et al., 2019), DistilBART (Shleifer & Rush, 2020).
-We also evaluate the recent, more efficient models, LLaMa3-8B (Touvron et al., 2023) and LLaMa3-70B (Taori et al., 2023).
-We assess the GPT-3.5 (i.e., ChatGPT), and the latest GPT-4 (OpenAI, 2023) by April 2023, using the OpenAI API (https://openai.com/api/) with temperature 0.
-We employed the most recent GPTO as well.
-Then, we assess C2P on a proposed synthetic dataset, similar to the one in (Jin et al., 2023a), but more complex and with less information provided on the premises.
-Then we test the reliability of C2P in different numbers of nodes.
+To test existing LLMs on the first synthetic data, we first include two BERT-based NLI models in the transformers library (Wolf et al., 2020): BART (Lewis et al., 2019), DistilBART (Shleifer & Rush, 2020).
+We evaluate LLaMA3-8B (Touvron et al., 2023) and LLaMA3-70B (Taori et al., 2023).
+We also evaluate the latest, more efficient models, LLaMA 3.1 8B and LLaMA 3.1 70B released in late September 2024.
+We assess the GPT-3.5 (i.e., ChatGPT), and the GPT-4 (Achiam et al., 2023), using the OpenAI API (https://openai.com/api/).
+We also used the latest GPT-4o model, released on October 2, 2024.
+Then, we evaluate the reliability of $\text{C}^{2}\text{P}$ on various queries, each with a different number of random variables.
 Due to the limitation of tokens for different versions of GPTs, we employed GP4-Turbo which has a 30000 maximum token limit.
-We selected 100 samples at random from the corr2cause dataset for our experiment, ensuring a balanced distribution with 50 "Yes" and 50 "No" answers.
-In the corr2cause dataset, the majority of responses are "No", as a result, they used the F1 score as the primary metric for accuracy.
-Our experiment design enhances both the realism and comprehensibility of the results.
+We also used $\text{C}^{2}\text{P}$ with LLaMA 3.1 70B as well.
+In all of the models, we set the temperature to 0.
+Note that various types of LLMs were tested in (Jin et al., 2023b), and the results showed that existing LLMs perform worse than random, completely random, or only slightly better than random in responding to such queries.
+From the models tested in (Jin et al., 2023b), we selected those with the highest accuracies to avoid duplicating the same results and to stay focused on the main objective of this study.
+According to Tab.
+LABEL:table01, the majority of responses in CORR2CAUSE dataset are "No" ($81.5\%$).
+Additionally, most samples involved in reasoning have six variables ($96.06\%$), a challenging task even for humans.
+As a result of these biases, the F1 score is used as the primary metric for evaluating accuracy in CORR2CAUSE simulations.
+We resolve these biases, by using 420 samples.
+First, to ensure a balanced distribution of samples, we select a population with 60 "Yes" and 60 "No" answers and an equal number of samples for each number of variables.
+This sample size is determined based on two main considerations.
+First, to ensure significant improvements in both step-by-step prompting and few-shot experiments, we applied the sample size formula for comparing two proportions as described by (Chow et al., 2017) and the minimum required sample size is 117.
+Additionally, to address the limited number of "Yes" responses (only 15 samples), we selected 30 samples for each variable number to minimize bias, resulting in a total of 120 samples.
+To ensure that the improvement of the results remains significant for the higher number of samples, we use 300 samples with 5 random variables with equal "Yes" and "No" answers.
+This experimental design enhances the realism and comprehensibility of the computational metrics, providing a more accurate reflection of the model’s performance.
+As a result, all four key metrics—F1 score, precision, recall, and accuracy—are practical, each offering a unique perspective on the model’s capabilities.
 These samples explore reasoning within three causal scenarios: direct cause-and-effect relationships (Fig.
@@
 $X_{1}$ and $X_{3}$ are causes of $X_{2}$ ($X_{2}$ is common effect of $X_{1}$ and $X_{3}$)
-To evaluate the proposed framework on natural stories, we utilized the state-of-the-art large language model, GPT-4, which excels at story generation.
-We crafted detailed instructions in the prompts and generated approximately 20 stories for our case study in fields such as healthcare and medicine, economics, education, social sciences, environmental science, and marketing, all emphasizing the importance of causality.
-This process is similar to the one presented in (Jin et al., 2023b), but our generated stories are more complex.
-Our approach can be tested using the examples of the Simpson’s Paradox discussed in (Pearl & Mackenzie, 2018).
+To evaluate the proposed framework for natural stories, we used GPT-4, which excels at story generation.
+We crafted instructions in the prompts and generated 30 stories for our case study in fields such as healthcare and medicine, economics, social sciences, environmental science, and marketing, all highlighting the importance of causality.
+This process is similar to the one presented in (Jin et al., 2023b).
+Our approach can be tested using the examples of Simpson’s paradox discussed in (Pearl & Mackenzie, 2018).
 However, since these examples are already included in the training data of current LLMs, the models simply repeat the correct answers based on that training data, similar to the parrot study in (Zečević et al., 2023).
 Consequently, our generated natural stories replicate these examples in a manner that the LLM cannot address within its existing training data.
-Additionally, we assess the coevolution of supermassive black holes (SMBHs) and their host galaxies using our proposed framework, replicating the results provided in (Pasquato et al., 2023).
-4.2 Evaluation of the C2P on synthetic dataset
-Results of the C2P on CORR2CAUSE dataset: In Table LABEL:table4, we show the performance of the more powerful LLMs in the cause-effect task reported in (Jin et al., 2023b) compared to the results of the C2P framework.
-Table 1: Main results: Comparison of using C2P frameworks compared to the existing LLMs
+We aimed to demonstrate how the symbolic expressions in the CORR2CAUSE dataset can affect the reasoning of our proposed framework.
+Finally, we assess the coevolution of supermassive black holes (SMBHs) and their host galaxies using our proposed framework, replicating the results from (Pasquato et al., 2023).
+The goal in (Pasquato et al., 2023) was to extract a PDAG of SMBHs based on numerical data and then infer causal reasoning questions from the graph.
+We use verbalized information on SMBHs to evaluate whether the LLM, enhanced with $\text{C}^{2}\text{P}$, can answer reasoning questions such as, "Does central density affect black hole mass?"
+It is important to note that Chain-of-Thought (CoT) prompting is implemented in most existing LLMs, enhancing their reasoning accuracy, as discussed in (Chung et al., 2024).
+Consequently, in our experiments, LLMs are prompted to think step by step when responding to reasoning questions, ensuring that the CoT mechanism is activated.
+4.3 Evaluation of the $\text{C}^{2}\text{P}$ on synthetic dataset
+Results of the $\text{C}^{2}\text{P}$ on CORR2CAUSE dataset: In Tab. 2, we show the performance of LLMs in the cause-effect task with and without employment of $\text{C}^{2}\text{P}$ framework.
+Table 2: Comparison of applying $\text{C}^{2}\text{P}$ frameworks in LLMs compared to the existing LLMs with CoT
 Models F1 Precision Recall Accuracy
@@
 LLaMA-Based Models
-LLaMa3-8B 48.97 51.06 47.05 51
-LLaMa3-70B 70.8 60.7 56.6 60
+LLaMA 3-8B 43.37 48.6 43.15 47.5
+LLaMA 3-70B 49.41 53.84 62.23 55.77
+LLaMA 3.1-8B 53.96 45.94 61.41 46.66
+LLaMA 3.1-70B 57.14 48.64 49.15 48.07
+$\text{C}^{2}\text{P}$ with LLaMA 3.1-70B 81.63 83.3 83.3 81.66
+$\text{C}^{2}\text{P}$ Few-shot learned LLaMA 3.1-70B 76 79.1 73.07 76.66
 GPT-Based Models
-GPT-3.5 48.8 55 44 54
-GPT-4 48.4 53.3 48 53
-GPT-4O 57.7 59.5 56 59
-C2P on GPT-4 Turbo 92.78 95.74 90 93
-Few-shot learned GPT-4 Turbo 81.06 82.97 79.24 81
-According to Table LABEL:table4, performing pure causal reasoning tasks presents a significant challenge for all existing LLMs.
-Notably, several models perform worse than a random guess, indicating a complete failure in this specific task of pure causal reason.
-Nevertheless, applying our proposed framework to GPT-4 Turbo has boosted at least 33% in the model’s ability to respond to causal inquiries, suggesting a substantial improvement over the existing capabilities of LLMs.
-It is important to note that the primary factor contributing to the discrepancy between the LLaMa and GPT results and those reported in (Jin et al., 2023b) stems from the distribution of the provided premise.
-This table indicates that when reasoning questions are posed to LLMs, the responses slightly tend towards the "No" answer.
-Robustness analysis: To assess the robustness of the proposed prompts, Table 2 shows the computed accuracy computed in each sub-task of implementing C2P on GPT-4 for different numbers of variables in the given premise.
-Table 2: Accuracy by Number of Variables and sub-tasks
+GPT-3.5 47.5 56.2 43.2 52.5
+GPT-4 Turbo 51.9 51.92 45 54.2
+GPT-4 50.2 54.1 47.1 55
+GPT-4O 59.86 61.46 58.33 60.83
+$\text{C}^{2}\text{P}$ with GPT-4 Turbo 91.72 93.47 89.2 91.66
+$\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo 81.6 79.21 77.21 79.28
+According to Tab.
+2, performing causal reasoning tasks remains a significant challenge for existing LLMs, even when prompted to think step by step, similar to (Wei et al., 2022b).
+These results show that even more updated models can sometimes perform worse than older ones in some metrics, for instance, LlaMa 3.1 acts worse than LlaMa 3.
+Based on the responses presented in the code repository for this study, model hallucination is one of the major factors contributing to this poor performance, aligning with the findings in Jin et al.
+(2023b).
+Tab.
+2 shows that by applying 5 consecutive tasks of $\text{C}^{2}\text{P}$ on GPT-4 Turbo and LlaMa 3.1, the reasoning accuracy improved by over $30.7\%$ and $25.9\%$, respectively, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on a synthetic benchmark dataset.
+Then, using few-shot learning of GPT-4 Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$, reasoning accuracy increased by over $20.05\%$ and $20.89\%$, respectively, with as few as ten examples, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on the same dataset.
+To ensure that the improvements are significant in both step-by-step prompting and few-shot experiments, we applied the sample size formula for comparing two proportions from (Chow et al., 2017).
+This formula indicates that even with a sample size of 117, the observed differences are statistically significant we assume a confidence level of $99\%$ and $80\%$ power.
+It is important to note that the primary factor contributing to the discrepancy between the results of LLaMA and GPT, and those reported in (Jin et al., 2023b), is the distribution of the provided premise.
+This table shows that when reasoning questions are posed to GPT models, their responses tend to slightly favor the "No" answer.
+Interestingly, LLaMA models slightly tend to answer "Yes" to reasoning questions more frequently.
+It is important to note that even when LLMs provide the correct final answer, the reasoning process that leads to that answer can still be incorrect, as illustrated in A.2, Tab.
+LABEL:tab:_wrong_and_right, for the results of GPT-4.
+This highlights the randomness of their responses more clearly which is also investigated and shown in Bao et al.
+(2024) as well.
+Other prompts that can execute the algorithm can be used in place of ours, as long as they perform the same subtasks.
+This supports the idea presented in the Lu et al.
+(2024).
+Furthermore, this becomes particularly evident in the few-shot learning process, where providing more examples has a significant impact; even more abstract prompts can achieve the same or even better results.
+Robustness and reliability analysis: To assess the robustness of the proposed prompts for each of the 5 subtasks, Tab.
+3 shows the computed accuracy computed in each subtask of implementing $\text{C}^{2}\text{P}$ (subsection 3.1) on GPT-4 Turbo for different numbers of variables in the given premise.
+Table 3: Accuracy by number of variables and subtasks
 Number of variables Accuracy
-First sub-task Second sub-task Third sub-task Fourth sub-task Fifth sub-task
-3 variables 100% 100% 100% 100% 99%
-4 variables 100% 100% 100% 98% 83.2%
-5 variables 100% 100% 100% 88.3% 75.2%
-6 variables 100% 100% 100% 79.1% 70.1%
-4.3 Evaluation of the C2P on natural stories
-To evaluate the proposed framework on natural stories, we utilize the same process presented in (Jin et al., 2023b) and generate premises.
-The result of applying the C2P framework, both step-by-step and few-shot learned, is provided in Table 3 and compared to the result of GPTs.
-This table clearly demonstrates that the existing method responds randomly to reasoning prompts.
-It is important to note that even when GPTs provide the correct final answer, the reasoning process leading to that answer is still incorrect.
-In contrast, GPT-4 Turbo, enhanced with step-by-step prompting of the C2P framework and few-shot learned C2P, boosts the accuracy of these models in similar scenarios involving natural stories.
-Table 3: Result of applying the C2P framework and GPTs, TP: True Positives, FP: False Positives, TN: true negatives, FN: False Negatives
+First subtask Second subtask Third subtask Fourth subtask Fifth subtask
+3 variables 100% 100% 100% 99.12% 98.7%
+4 variables 100% 100% 100% 97.5% 84.1%
+5 variables 100% 100% 100% 87.5% 75%
+6 variables 100% 100% 100% 78.3% 70%
+The sample size formula for comparing two proportions indicates that fewer than 120 samples are sufficient to demonstrate a significant improvement in reasoning performance when using $\text{C}^{2}\text{P}$.
+Figure 3 shows results for datasets with sample sizes ranging from 30 to 300.
+This figure demonstrates that LLMs, such as GPT-4 Turbo, consistently achieve higher performance across all four metrics when $\text{C}^{2}\text{P}$ is applied, compared to scenarios where $\text{C}^{2}\text{P}$ is not used.
+Figure 3: The accuracy metrics for different numbers of samples with 3 different methods.
+According to this figure, the $\text{C}^{2}\text{P}$ framework enhances all metrics across varying sample sizes.
+4.4 Evaluation of the $\text{C}^{2}\text{P}$ on natural stories
+The results of applying the $\text{C}^{2}\text{P}$ framework, both step-by-step and few-shot learned for GPT-4 Turbo and LLaMA 3.1, are provided in Tab.
+4 and compared to the results of GPTs and LLaMA 3.1.
+This table, along with the highlighted sections, clearly demonstrates that the existing methods respond randomly to reasoning prompts.
+In contrast, GPT-4 Turbo and LLaMA 3.1 with step-by-step prompting of the $\text{C}^{2}\text{P}$ framework and few-shot learned $\text{C}^{2}\text{P}$, significantly improve accuracy in scenarios involving natural stories.
+Table 4: Comparison of applying $\text{C}^{2}\text{P}$ frameworks with existing LLMs with CoT in responding to queries within natural stories, TP: True Positives, FP: False Positives, TN: true negatives, FN: False Negatives.
 Models TP FP TN FN
-GPTs
-GPT-4.5 2 2 3 3
-GPT-4 2 2 3 3
-GPT-4o 3 2 3 3
-Featured GPT
-C2P on GPT-4 Turbo 5 0 5 0
-Few-shot learned GPT-4 Turbo 4 0 5 1
-Additionally, Table 4 provides an example of such natural stories and demonstrates the result of applying a few-shot learned GP4-Turbo with C2P to it.
+LLMs without $\text{C}^{2}\text{P}$
+GPT-3.5 5 8 7 10
+GPT-4 Turbo 6 7 8 9
+GPT-4 7 8 7 8
+GPT-4O 8 6 9 7
+LLaMA 3-70B 9 9 6 6
+LLaMA 3.1-70B 9 10 5 6
+Featured LLMs with $\text{C}^{2}\text{P}$
+$\text{C}^{2}\text{P}$ on LLaMA 3.1-70B 11 3 12 4
+$\text{C}^{2}\text{P}$ Few-shot learned LLaMA 3.1-70B 10 3 12 5
+$\text{C}^{2}\text{P}$ on GPT-4 Turbo 13 1 14 2
+$\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo 11 3 12 4
+Additionally, Tab.
+5 provides an example of such natural stories and demonstrates the result of applying few-shot learned GP4-Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$ to them.
 This showcases how GPT models (including versions 3.5, 4, 4 Turbo, and 4O) fail to respond correctly.
-It is important to note that the few-shot learning process of GPT-4 Turbo was performed using only six examples from the CORR2CAUSE dataset, without including any natural story examples.
-Note that there are no natural stories employed in the few-shot learning process.
-Table 4: Example of a natural story and the result of few-shot learned GPT-4-Turbo with C2P
-Example of true story Results of few-shot learned GPT-4 Turbo with C2P
+It is important to note that the few-shot learning process of GPT-4 Turbo and LLaMA 3.1 was performed using only ten examples from the CORR2CAUSE dataset, without including any natural story examples.
+Table 5: Example of a natural story and the result of few-shot learned GPT-4-Turbo with $\text{C}^{2}\text{P}$
+Example of true story Results of few-shot learned GPT-4 Turbo with $\text{C}^{2}\text{P}$
 Premise: eating junk food (A), obesity (C), and watching television (B) have relations with each other.
@@
 However, eating junk food and watching television are independent from each other.
-Hypothesis: Eating junk food directly affects obesity Step 1: Number of random variable:3 Names of random variable: A, B, C
-Step 2: All of Statistical Relations:{ Dependencies: [[A, C], [B, C]], Unconditional Independencies: [[A, B]], Conditional Independencies: []}
-Step 3: A: {A: 0, B: 1, C: 1}, B: {A: 1, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
-Step 4: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
-Step 5: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 1, B: 1, C: 0}
+Hypothesis: Eating junk food directly affects obesity Step 1: Number of random variable:3
+Names of random variable: A, B, C
+Step 2: All of Statistical Relations:{
+Dependencies: [[A, C], [B, C]],
+Unconditional Independencies: [[A, B]],
+Conditional Independencies: []}
+Step 3: A: {A: 0, B: 1, C: 1},
+B: {A: 1, B: 0, C: 1},
+C: {A: 1, B: 1, C: 0}
+Step 4: A: {A: 0, B: 0, C: 1},
+B: {A: 0, B: 0, C: 1},
+C: {A: 1, B: 1, C: 0}
+Step 5: A: {A: 0, B: 0, C: 1},
+B: {A: 0, B: 0, C: 1},
+C: {A: 1, B: 1, C: 0}
 Step 6: C: [A, B]
 Step 7: C: [A, B]
-Step 8: A: {A: 0, B: 0, C: 1}, B: {A: 0, B: 0, C: 1}, C: {A: 0, B: 0, C: 0}
-Step 9: Checking matrix[A][D] = 1 and matrix[D][A] = 0. According to rule 2, this suggests A is a direct cause of D, or D is a direct effect of A.
-4.4 Evaluation of the C2P framework on the (co)evolution of supermassive black holes and their host galaxies
-In astrophysics, experiments are impossible.
+Step 8: A: {A: 0, B: 0, C: 1},
+B: {A: 0, B: 0, C: 1},
+C: {A: 0, B: 0, C: 0}
+Step 9: Checking matrix[A][C] = 1 and matrix[C][A] = 0. According to rule 2, this suggests A is a direct cause of C, or C is a direct effect of A.
+4.5 Evaluation of the $\text{C}^{2}\text{P}$ framework on the (co)evolution of supermassive black holes and their host galaxies
+In astrophysics, interventional experiments are impossible.
 Thus, causal reasoning must rely exclusively on observational data.
-Similar to the study in (Pasquato et al., 2023), we (co)evolute supermassive black holes (SMBHs) and their host galaxies.
-We implement C2P on GPT-4 Turbo to a comprehensive catalog of galaxy properties to obtain a PDAG and then reason the causal questions.
-Fig. 3 shows the results of each sub-tasks of implementing C2P
-Figure 3: Prompts (Q) and results (A) of sub-tasks application of the C2P framework to real-world complex scenarios and steps of sub-task 3 for the given premise.
-According to the extracted PDAG, many questions such as the causal given can be answered which is too complicated for the existing LLMs to answer.
-It is important to note that by implementing C2P, not only many causal questions can be answered with this approach, but it is easy to show which causal questions can not be answered based on the given premise and which more information has to be given to be able to causally reason.
-More importantly, it is worth noting that this is not the limitation of the C2P framework, any rational agent needs such information to answer such causal questions.
-5 Discussion, Limitations and Future Work
-Causal reasoning and identifiability: Having a PDAG of an SCM of a causal mechanism is the most effective way to achieve this with observational data.
-Based on this, two main questions arise: Given a PDAG, under what conditions can we make causal reasoning?
-This involves determining the necessary assumptions and data required to address a causal question—this issue is known as identification.
+Pasquato et al.
+(2023) computationally studied the coevolution of supermassive black holes based on numerical data, and a causal graph for the underlying mechanism is extracted.
+Then, the causal hypothesis, "central density affects black hole mass", is answered.
+Similar to their study, we (co)evolute supermassive black holes (SMBHs) and their host galaxies based on the verbalized information in (Pasquato et al., 2023).
+We GPT-4 Turbo with $\text{C}^{2}\text{P}$ to process the verbalized information provided in subsection 4.1, obtain a PDAG, and then reason about the causal questions.
+Fig. 4 illustrates the results of the subtasks of $\text{C}^{2}\text{P}$ on SMBHs data presented.
+Figure 4: Prompts (Q) and results (A) of subtasks application of the $\text{C}^{2}\text{P}$ framework to real-world complex scenarios and steps of subtask 3 for the given premise.
+The first four subtasks aim to extract the PDAG from the given data.
+Then, the hypothesis that central density affects black hole mass is evaluated in subtask 5 and answered.
+According to the extracted PDAG, many other reasoning questions questions can be answered that are too complicated for the existing LLMs to answer.
+It is important to note that by implementing $\text{C}^{2}\text{P}$, not only many causal questions can be answered with this approach, but it is easy to show which causal questions can not be answered based on the given premise and which more information has to be given to be able to causally reason.
+The discussion on which questions can and cannot be answered based on PDAG is presented in detail in studies such as (Hernán & Robins, 2006; Hauser & Bühlmann, 2012; Perkovic, 2020).
+Importantly, the inability to answer certain causal reasoning questions based on the PDAG, or its corresponding adjacency matrix, does not reflect a limitation of the $\text{C}^{2}\text{P}$ framework.
+Rather, any rational agent needs additional information to effectively address all causal inquiries.
+5 Discussion, Challenges and Future Work
+Practical insights on simulations and results of $\text{C}^{2}\text{P}$: Each part of our study has a specific aim.
+The goal of the simulations, where we applied the 9 steps of $\text{C}^{2}\text{P}$ step by step, is to show what the expected output would be and how it leads to causal reasoning.
+However, our investigations demonstrate sensitivity to the prompts.
+In few-shot learning, where examples of expected results are provided for each prompt, the sensitivity to prompts was not significant; instead, the results were primarily dependent on the number of examples.
+Interestingly, while the prompt is lengthy —due to the fact that all steps are given to LLMs as a single prompt —- the results remained highly accurate with as few as ten examples, as reported in Tab.s LABEL:table4 and 5.
+By comparing the prompts in Appendices A.3 and A.4, it is clear that the prompts in the few-shot learning section do not need to be as detailed and are more abstract than those used in the step-by-step application of $\text{C}^{2}\text{P}$.
+Additionally, it is important to note that the results for natural stories using few-shot learned $\text{C}^{2}\text{P}$ were highly accurate (see Table 5), even though the examples in the few-shot learning process were synthetic “cause-effect" examples from the CORR2CAUSE dataset rather than real story examples.
+The comparison of results between $\text{C}^{2}\text{P}$ and CoT provides new insights into how LLMs aim to reason and why $\text{C}^{2}\text{P}$ and similar frameworks can guide a model to "think" step by step.
+As demonstrated in the examples provided in the code repository—under "Sample Responses of LlaMa 3.1 with CoT" for LlaMa 3.1 and "Sample Responses of GPT-4 Turbo with CoT" for GPT-4 Turbo— both models attempt to break tasks into multiple subtasks when they are asked to think step by step (sometimes with more than nine consecutive subtasks).
+However, they still lack rationality in how they structure these subtasks, as it is discussed in Wei et al.
+(2022a).
+In other words, they struggle to design the necessary subtasks for reasoning, a task that $\text{C}^{2}\text{P}$ successfully performs to enhance reasoning.
+This challenge is also highlighted in Wei et al.
+(2022b), where CoT is introduced, and its logical inconsistencies and poor step alignment are discussed as key limitations of the approach.
+Causal reasoning and identifiability: Extracting a PDAG using only observational data is a key step in learning the true underlying causal mechanism.
+Based on the extracted PDAG, two main questions arise: Given a PDAG, under what conditions can we make causal reasoning?
+This involves determining the necessary assumptions and data required to address a causal question.
+This issue is known as identifiability.
 It is generally possible to answer some cause-effect questions based on the PDAG, represented by a causal graph, and lower-level data.
@@
 As a result, the second question is, what else is needed to answer all the causal questions?
-In such cases, interventional data is necessary to explicitly extract the direction of causal relations in a mechanism.
+In such cases, interventional data is necessary to explicitly extract all the directions of causal relations in a mechanism.
 To do so, while a naive baseline approach would require $O(n^{2})$ interventions, various methods have been proposed, such as those in (Kocaoglu et al., 2017; Choo & Shiragur, 2023; von Kügelgen et al., 2024; Squires & Uhler, 2023), for cases in different situations.
@@
 PDAGs, with their partial specification, might not support detailed predictions under interventions without resolving the ambiguities in causal direction.
+Next Steps for Reasoning in LLMs: The human brain possesses an intrinsic drive to understand causality.
+Whether driven by curiosity or the pursuit of specific goals, we continuously seek to understand why events occur and how they are interconnected.
 Causal reasoning is a broad and complex task in AI and LLMs.
-While current machine learning methods struggle with extracting causal structures and subsequently reasoning causally, this problem is even more intricate in the context of LLMs.
-The primary reason for this complexity is the inherent structure of LLMs and their reliance on the concept of attention.
-To enable LLMs to reason causally, the first step is to identify the causal question.
-Recently, a comprehensive database was compiled, and an AI model was trained to identify causal questions (Ceraolo et al., 2024).
-This model can be used prior to applying our framework to determine whether a causal reasoning process has to be activated or not.
+While current machine learning methods find it difficult to extract causal structures and subsequently reason causally, this problem can be even more intricate within the context of LLMs.
+The primary reason for this complexity is the distinction between association and cause-effect relationships.
+The inherent structure of LLMs relies on the attention mechanism presented in (Vaswani et al., 2017), which is akin to what Pearl refers to as "association" in the ladder of causality (Pearl & Mackenzie, 2018).
+As a result, advancing up the ladder of causality is essential for enabling true reasoning.
+To enable LLMs to address this fundamental quest for causality, they must be capable of sequentially performing two main tasks.
+As demonstrated in our study, $\text{C}^{2}\text{P}$ has significantly enhanced the reasoning capabilities of LLMs on the provided datasets, which can be the main task in the reasoning process.
+However, the first task is to be able to understand the causal questions, which involves formalizing the definition of causal questions and establishing a taxonomy for finer-grained classification.
+The initial effort on this is presented in Ceraolo et al.
+(2024).
+Then, $\text{C}^{2}\text{P}$ can be employed.
+However, the first subtask of this framework is to extract random variables, which requires the models to understand the definition of a random variable to ensure that the variables in the reasoning process are correctly identified.
+Failure in this process could undermine our proposed framework.
+Another important concept in this process is extracting the dependencies provided in the premise.
+In the datasets used for our experiments, these dependencies are explicitly stated in the given premise; however, in reality, they are often implicit.
+As a result, LLMs should be able to accurately extract these dependencies.
 Additionally, more comprehensive examples and scenarios need to be generated to aid in the learning process of an LLM.
-Due to the vast variety of causal scenarios and the token limitations of current models, our approach must be integrated into the learning or fine-tuning process of an LLM.
-This ensures that the model can effectively learn and apply causal reasoning within its capabilities.
-Scaling Instruction-Finetuned Language Models: LLMs primarily function through pattern recognition without the capability to recognize causal relationships, as indicated by prior research .
-This limitation highlights the importance of implementing these models with causal analysis techniques to enhance their causal reasoning ability.
-Previous attempts to boost LLMs’ causal reasoning involved using external tools or manually crafted prompts, which, although insightful, are impractical for easy integration.
-Our study introduced an algorithm that leverages LLMs’ inherent abilities to foster causal reasoning autonomously.
-Our findings demonstrate a marked increase in the accuracy of LLMs’ causal understanding, even in scenarios involving complex causal interactions.
-Additionally, incorporating a chain of thought into LLMs’ training has been proven to enhance performance across various tasks (Chung et al., 2024).
-Based on the extracted results on the benchmark and provided datasets in our study, and the results of the few-shot learning process, we strongly hypothesize that our approach could similarly benefit the training of language models, particularly in boosting their causal reasoning, even in smaller models that would lead to a revelation in LLMs.
-This enhancement could be realized using a carefully selected set of examples or synthetic samples created by LLMs, following the guidelines outlined in our algorithm.
-Future works As demonstrated in the results section, C2P has significantly enhanced the reasoning capabilities of LLMs.
-We have shown that extracting the adjacency matrix of the variables in a given premise can function as the DAG in Pearl’s works, leading to improved reasoning ability.
-Our experimental results of few-shot learned LLMs indicate a significant increase in accuracy for tasks where users inquire about direct cause-and-effect relationships or common cause questions.
-However, the models still struggle with indirect cause-and-effect questions.
-Pearl demonstrated that the concepts of a DAG and d-separation complement each other in causal reasoning tasks.
-To address the low accuracy in answering indirect causal reasoning questions, more sub-tasks can be introduced to LLMs to perform d-separation in a more structured manner.
-As mentioned, due to the token limitation of LLMs, we were able to use only six or even fewer examples in the few-shot process of GPT-4-Turbo.
-As a result, our proposed framework can be seamlessly integrated into the fine-tuning or learning process of an LLM by introducing more extensive examples with more complex scenarios.
-Given the accuracy achieved in few-shot learned LLMs, the fine-tuned or trained LLMs with our approach can achieve highly accurate results, effectively addressing the causal reasoning bottleneck of LLMs.
+By overcoming these challenges, the integration of $\text{C}^{2}\text{P}$ with LLMs can provide these models with causal reasoning capabilities, similar to the transformative impact of "Chain-of-Thought" (Wei et al., 2022b), as highlighted by Chung et al.
+(2024).
+It is also important to note that while our experimental results with few-shot learned LLMs show a significant increase in accuracy for tasks involving direct cause-and-effect relationships or common cause questions, the models still struggle with indirect cause-and-effect questions.
+Moreover, the accuracy of our framework decreases as the number of random variables increases.
+As Pearl demonstrated that DAGs and d-separation are complementary in causal reasoning tasks, more structured subtasks can be introduced to improve performance in answering indirect causal questions and handling more complex scenarios with numerous variables.
+These subtasks would help guide LLMs in performing reasoning tasks more effectively.
+The key benefit of few-shot learning is that it demonstrates that pre-trained models can perform causal reasoning without needing to be retrained from scratch or necessarily fine-tuned, making it more cost-effective and compute-efficient.
+This leads to reduced costs and less demanding infrastructure requirements.
+Therefore, a model fully trained or fine-tuned with $\text{C}^{2}\text{P}$ would probably perform even better than few-shot learned ones, as it can be trained or fine-tuned on thousands of examples instead of the ten examples used during few-shot learning.
+This highlights how integrating $\text{C}^{2}\text{P}$ during the training and fine-tuning of LLMs can revolutionize existing models.
+This is also discussed in Chung et al.
+(2024) as the “scaling law”.
+Additionally, Studies like Kaplan et al., 2020 showed that language models benefit from increased data and computational resources.
+More data during fine-tuning allows the model to capture better representations, leading to improved performance in various tasks.
+This is one of the core principles behind why larger datasets, would likely boost accuracy compared to just 10 examples.
+In Brown et al.
+(2020) with GPT-3, the authors highlight how performance increases as the model is trained on more examples.
+In few-shot learning, the model can generalize from a small set of examples, but fine-tuning with larger datasets usually leads to more substantial improvements.
+Devlin (2018) also demonstrates that fine-tuning with more task-specific examples improves the model’s performance.
+As a result, by addressing the mentioned challenges, employing $\text{C}^{2}\text{P}$ in the fine-tuning process is expected to improve the reasoning accuracy of the models.
 Appendix A Appendix
-A.1 Prompts for C2P frameworks
-The proposed prompts for step-by-step proceeding with C2P are as follows:
+A.1 do-calculus rules
+The three main rules of the do-calculus are as follows:
+  1. 1.
+    Insertion/deletion of observations: If a variable $X_{2}$ is independent of $X_{1}$ given a set of other variables $X_{3}$, then the conditional distribution of $X_{2}$ given $do(X_{1})$ and $X_{3}$ is the same as the conditional distribution of $X_{2}$ given $X_{3}$ alone.
+This rule can be applied when $X_{1}$ and $X_{2}$ are d-separated by $X_{3}$ in the graph modified by removing all arrows pointing into $X_{1}$.
+  2. 2.
+    Action/observation exchange: If $X_{2}$ is independent of $X_{1}$ given $X_{3}$ and another set $W$ where $X_{1}$ is intervened, then the graph can be modified to reflect this exchange.
+  3. 3.
+    Insertion/deletion of actions: This rule states that if intervening on $X_{1}$ does not affect $X_{2}$ after accounting for $X_{3}$, the intervention can be ignored in the expression.
+A.2 Example of comparison in responding to causal queries
+For the given example in Subsection 4.1, Tab.
+LABEL:tab:_wrong_and_right demonstrates that while GPT-4 arrives at the correct final answer, the reasoning process is flawed, providing further evidence of the random behavior of existing models in causal reasoning.
+Table 6: An example of the correct final answer with GPT-4 Turbo while the reasoning process leading is incorrect and results of $\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo
+GPT-4 Turbo with CoT $\text{C}^{2}\text{P}$ Few-shot learned GPT-4 Turbo
+The hypothesis that A directly affects C could be possible.
+The premise states that A correlates with C, which means there is a relationship between them.
+However, correlation does not necessarily imply causation.
+It could be that A affects C, C affects A, or that both A and C are affected by a separate, unmentioned variable.
+However, since the premise does not provide any information contradicting the hypothesis, it is possible that A directly affects C.
+Final Answer: "Yes" According to the adjacency matrix, matrix["A"]["C"] = 1 and matrix["C"]["A"] = 0 indicates that A directly cause C.
+Therefore, the hypothesis that A causes C or C is effect of A is correct.
+Final Answer: "Yes"
+A.3 Prompts for $\text{C}^{2}\text{P}$ frameworks
+The proposed steps (prompts) for step-by-step proceeding with 5 main subtasks of $\text{C}^{2}\text{P}$ are as follows:
   1. 1.
     Prompting for extracting the random variables from the provided data.
-    1. (i)
-      Please give me the number of random variables in the given premise
-    2. (ii)
-      Write the names of all random variables.
+    • Step 1: Please give the number of random variables in the given premise and write the names of all random variables.
   2. 2.
     Prompting to extract all the cause and effect relations along with all conditional and unconditional relations among the random variables specifically mentioned in the given premise
-    • If 2 random variables, for instance, $R_{1}$ and $R_{2}$, are independent, write it in this form: "$R_{1}$ is independent of $R_{2}$".
-If there exist 2 random variables, for instance, $R_{1}$ and $R_{2}$, are conditionally independent given a third random variable, for instance, $R_{3}$, write it in this form: "$R_{1}$ and $R_{2}$ are independent given $R_{3}$".
-If two random variables, for instance, $R_{1}$ and $R_{2}$, are specially mentioned to have cause and effect relation, write it in this form: "$R_{1}$ is the cause of $R_{2}$".
+    • Step 2: If 2 random variables, for instance, $R_{1}$ and $R_{2}$, are independent, write it in this form: $R_{1}$ is independent of $R_{2}$".
+If there exist 2 random variables, for instance, $R_{1}$ and $R_{2}$, are conditionally independent given a third random variable, for instance, $R_{3}$, write it in this form: $R_{1}$ and $R_{2}$ are independent given $R_{3}$".
+If two random variables, for instance, $R_{1}$ and $R_{2}$, are specially mentioned to have cause and effect relation, write it in this form: $R_{1}$ is the cause of $R_{2}$".
   3. 3.
     Prompting to create an adjacency matrix where all elements are 1, except for the diagonal elements and the elements corresponding to the cause-and-effect relationships specifically mentioned in the given premise.
-    • In this phase, each random variable is treated as a node within a fully connected undirected graph.
-Then, for each pair, for instance, $R_{1}$ and $R_{2}$, presented in the form: "$R_{1}$ is the cause of $R_{2}$", set the element in ["$R_{2}$", "$R_{1}$"] in the adjacency matrix to 0.
+    • Step 3: In this phase, each random variable is treated as a node within a fully connected undirected graph.
+Then, for each pair, for instance, $R_{1}$ and $R_{2}$, presented in the form: $R_{1}$ is the cause of $R_{2}$", set the element in $R_{2}$", $R_{1}$"] in the adjacency matrix to 0.
   4. 4.
     Prompting of the conditional and unconditional independency valuation and identifying the colliders, step by step, to extract the causal PDAG.
-    1. (i)
-      Update the adjacency matrix based on the specified unconditional independencies between random variables.
+    • Step 4: Update the adjacency matrix based on the specified unconditional independencies between random variables.
 Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix.
@@
 Instructions: - For each pair of variables listed as independent, set their corresponding entries in the adjacency matrix to 0.
-    2. (ii)
-      Update the adjacency matrix based on the specified conditional independencies between random variables.
+    • Step 5: Update the adjacency matrix based on the specified conditional independencies between random variables.
 Each pair of variables that is declared independent should have their corresponding value set to zero in the adjacency matrix.
@@
 Instructions: - For each pair of variables listed as independent given other variable(s), set their corresponding entries in the adjacency matrix to 0.
-    3. (iii)
-      Task: Given an initial adjacency matrix, follow these steps: Step 1: Identify all rows (key values) in the matrix where there are two or more than two columns with the value "1" in them.
+    • Step 6: Task: Given an initial adjacency matrix, follow these steps: Step 1: Identify all rows (key values) in the matrix where there are two or more than two columns with the value "1" in them.
 For each identified row, find all pairs of different columns where the values are "1".Ensure to exclude rows that do not contain any pairs from the results.
 Step 2: Display these pairs, "All Pairs", where each row name is key, and the value is a list of column names that are identified in Step 1.
-    4. (iv)
-      Given the "All Pairs" and the list of independencies, follow these instructions step by step: Instruction: For each key in "All Pairs", delete all the pairs that are not mentioned as independent in the "independencies" list and return other with all their values.
+    • Step 7: Given the "All Pairs" and the list of independencies, follow these instructions step by step: Instruction: For each key in "All Pairs", delete all the pairs that are not mentioned as independent in the "independencies" list and return other with all their values.
 The "All Pairs" contains pairs of elements associated with each key.
@@
 The list of independencies provides information about which pairs are independent of each other.
-    5. (v)
-      Given the initial adjacency matrix represented and the "All Pairs" list, for each key-value pair ("$R$") in "All Pairs", modify the initial adjacency matrix as follows: 1- Set the value in the "$C_{1}$" row and "$R$" column to 0: ("$C_{1}$", "$R$") = 0.
-2- Set the value in the "$C_{2}$" row and "$R$" column to 0: ("$C_{2}$", "$R$") = 0.
+    • Step 8: Given the initial adjacency matrix represented and the "All Pairs" list, for each key-value pair ("RR") in "All Pairs", modify the initial adjacency matrix as follows: 1- Set the value in the $C_{1}$" row and "RR" column to 0: ($C_{1}$", "RR") = 0.
+2- Set the value in the $C_{2}$" row and "RR" column to 0: ($C_{2}$", "RR") = 0.
   5. 5.
     Prompting for cause-and-effect questions or hypotheses
-    • To extract and understand causal relations in the adjacency matrix: For each specified variable "$R$" and "$C$", for instance, that are listed in the adjacency matrix: - If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 1, then the causal direction between "$R$" and "$C$" is undetermined.
-- If matrix entry at ["$R$", "$C$"] = 1 and ["$C$", "$R$"] = 0, then "$R$" is a direct cause of "$C$" or "$C$" is a direct effect of "$R$".
-- If matrix entry at ["$R$", "$C$"] = 0 and ["$C$", "$R$"] = 1, then "$C$" is a direct cause of "$R$" or "$R$" is a direct effect of "$C$".
+    • Step 9: To extract and understand causal relations in the adjacency matrix: For each specified variable "RR" and "CC", for instance, that are listed in the adjacency matrix: - If matrix entry at [ "RR", "CC"] = 1 and ["CC", "RR"] = 1, then the causal direction between "RR" and "CC" is undetermined.
+- If matrix entry at ["RR", "CC"] = 1 and ["CC", "RR"] = 0, then "RR" is a direct cause of "CC" or "CC" is a direct effect of "RR".
+- If matrix entry at ["RR", "CC"] = 0 and ["CC", "RR"] = 1, then "CC" is a direct cause of "RR" or "RR" is a direct effect of "CC".
 If two variables directly affect a third variable, the first two variables are common causes, and the third variable is a collider.
-Evaluate the hypothesis based on the given partially directed acyclic graph presented as an adjacency matrix with the given Instruction.
-A.2 Prompts for few-shot learning of C2P
-An example of given prompts for the few-shot learning process of C2P is as follows:
+Evaluate the hypothesis based on the given partially presented as an adjacency matrix with the given Instruction.
+A.4 Prompts for few-shot learning of $\text{C}^{2}\text{P}$
+An example of given prompts for the few-shot learning process of $\text{C}^{2}\text{P}$ is as follows:
 Premise: Suppose there is a closed system of 5 variables, A, B, C, D, and E.